The achicago BibTeX style

A Chicago Manual BibTeX style

           Matt Swift <swift@alum.mit.edu>

Version: 1 Date: 2001/08/31

Documentation revision: 2001/08/31

To do: explain conventions; put some in compsci

Abstract

Chicago Manual of Style bibliography style BibTEX macros.

Contents

I Discussion 3

1 Pros and cons of this style 3

2 User documentation 3

II Implementation 6

3 Version control 6

4 Implementation 6

4.1 Define fields and variables . . . . . . . . . . . . . . . . . . . . . . . 6 4.1.1 Global variables . . . . . . . . . . . . . . . . . . . . . . . . 6 4.1.2 Bibfields . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

4.2 Define macros . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 4.3 Debugging functions . . . . . . . . . . . . . . . . . . . . . . . . . . 8 4.4 Initialization functions . . . . . . . . . . . . . . . . . . . . . . . . . 9 4.5 Boolean logic . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 4.6 Constant functions . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 4.7 Output functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 4.8 Simple auxiliary functions . . . . . . . . . . . . . . . . . . . . . . . 13 4.9 Queue functions START FIXING DOX HERE . . . . . . . . . . . 16 4.10 Format functions dox . . . . . . . . . . . . . . . . . . . . . . . . . . 17 4.11 Format names functions . . . . . . . . . . . . . . . . . . . . . . . . 18 4.12 Make functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 4.13 Label and principal field functions . . . . . . . . . . . . . . . . . . 31 4.14 Begin and end entries . . . . . . . . . . . . . . . . . . . . . . . . . 39 4.15 Entry types . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39

4.15.1 Article . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 4.15.2 Book . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 4.15.3 Inbook . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 4.15.4 Incollection . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 4.15.5 Majorthesis . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 4.15.6 Manual . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 4.15.7 Booklet . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 4.15.8 Inproceedings . . . . . . . . . . . . . . . . . . . . . . . . . . 48 4.15.9 Proceedings . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 4.15.10 Misc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 4.15.11 Techreport . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 4.15.12 Unpublished . . . . . . . . . . . . . . . . . . . . . . . . . . 50 4.15.13 Default . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51

4.16 Sorting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51

5 Disambiguating citations and sorting the bibliography 53 5.1 Actions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55

Part I

Discussion

Warning: The dox are spotty for this bibstyle—sorry! This is a pretty stable style, however. I’ve included almost my entire test cases. If you’re puzzled, you may compare frankenstein.bib with what you see in the References section at the end of this document. Or please write me.

Reference list ordering: alphabetical by author or whatever passes for author in the absence of one.

1 Pros and cons of this style

See a discussion in the documentation for the achicago package.

2 User documentation

No bibstyle is going to elgantly handle everything while using them in a way that’s at all compatible with the majority of existing bibstyles out there. And it’s not clear that such a bibstyle would be worth the effort to write and learn to use properly. The time would be better spent, I think, writing some macros to make inserting hand-done exceptions seamlessly into the bibliography, so that the whole process could remain clean and automatic.

For the sceptical, and for your entertainment, an example. I wanted to a cite a certain paper by Freud in my dissertation. The cited paper has a title, date of first publication, and chapter number; I read this paper in a volume of Freud’s papers with a title, volume editor, volume date, publisher, reprint date, and reprint publisher; this volume has a number in a series of publications (The International Psycho-Analytic Library) with a series editor; and also this volume has a volume number within a multivolume work with its own title (Collected Papers) and (supervising) translator. Readers honestly don’t need all that information, but each one of those 15 pieces of information might be highly important for another citation, for which only two or three pieces of information were relevant or available. No one needs to write a bibstyle that handles it all, so at several places, the fields are assumed to be used for either one thing or another. For example, ‘’ series either contains the title of the inclusive multivolume work or the series. When, as rarely happens, a volume is part of both at once, you have to give up one or the other.

Warning: This is a complex bibstyle and the number of potential cases is very high. Please check your output, and write me with cases that are not set according to Chicago Manual percepts. Please refer to Chicago Manual whenever possible.

To do: star the new fields.

entry type  description

address Address of publisher of cited edition, whether a reprint or

not.

annotation A complete paragraph or more.

Table 1: Entry types.

entry type  description

author bookauthor Book’s author whenever ‘’ author is referring to the author

of a part of the book. Usually an ‘inbook’ or “incollection” entry refers to a book with an editor but no author, but in the case of introductions etc. you have an author for each.

booktitle Book’s title whenever ‘’ title is referring to a part of the

book.

chapter The ‘’ title names the ‘’ chapterth part of the book. One

expects to have nonempty ‘’ booktitle. See also ‘’ type.

edition Edition information to be given after the book’s title, which is ‘’ booktitle if that is nonempty and ‘’ title other-

              wise. Example: 2d~ed.,2~vols.in1~book

editor NOTE all f* and o* fields are only going to apply to ‘book’, “incollection”, ‘inbook’ entries. FIX: check – booklet? etc.

faddress Address of publisher of first-published edition. Not used

unless ‘’ fyear is nonempty.

flanguage Language of first-published edition. Not used unless ‘’

translator or ‘’ fyear is nonempty.

fpublisher Publisher of first-published edition. Not used unless ‘’

fyear is nonempty.

ftitle Title of first-published edition (probably in ‘’ flanguage).

Not used unless ‘’ fyear is nonempty.

fyear Publication date of first-published edition. howpublished illustrator For ‘book’ entries only. When the author or editor illus-

trates, use a construction like {theeditor}andothers.

institution For thesis entries only. FIX check journal For ‘article’ entries only. key often default with empty author, editor, organization month For ‘journal’ entries, also season, e.g., Winter note Begins its own block, so capitalize. Concluding puctuation

is supplied automatically if not given in the field.

number For ‘journal’ entry, number of the journal (if no “” volume) or number within the volume. For “book” entry, number of the book in the ‘’ series. Makes sense only when ‘’ series is a series rather than the title of a multivolume work.

oaddress For a reprint, address of ORIGINAL publisher. opublisher For a reprint, publisher of ORIGINAL edition. A

nonempty value implies the cited edition IS a reprint (and so must have ‘’ oyear).

organization oyear For a reprint, publication date of ORIGINAL edition A

nonempty value implies the cited edition IS a reprint (and so OUGHT TO have ‘’ opublisher). The value will almost always differ from ‘’ year.

pages publisher Publisher of the cited edition.

Table 1: Entry types.

entry type  description

school series Series to which ‘’ title belongs, or title of a multivolume

work where the volumes have their own titles given in ‘’ booktitle or ‘’ title.

seriesedition Edition information to be given after ‘’ series. Makes sense only when ‘’ series is the title of a multivolume work with titled volumes. In the entry for the multivolume work, that is, when ‘’ series and ‘’ title are identical, use ‘’ seriesedition not ‘’ edition. See also ‘’ edition. Consequence of disobeying last is when references are less than min-crossrefs (i.e., no separate entry for the multivolume work) you get the edition information for the multivolume work appearing after the title of the component volume. Note: you can’t really get away with putting a series editor here. I think you’d have to use the ‘’ note field for that; or we could make yet another field.

title translator For ‘book’ entries only, the translator(s). See also “” flan-

guage. When the author or editor translates, use a con-

struction like {theeditor}andothers.

type For ‘inbook’ entries, the value overrides Chapter. When “” type is empty you get Chapter~<chapter>of. FIX: CHECK: For *-thesis entries, gives thesis-type word, e.g.

Master'sThesis.

volume For ‘journal’ entry, it’s volume. For “book” entry, its number in the ‘’ series, or its volume number in a multivolume book.

year Year of publication of cited edition, whether a reprint or

not.

yearcomp Year work composed. For ‘book’ entry only (so far). Table 1: Entry types.

Part II

Implementation

3 Version control

Version information:

 1 ^^A Keep the /^%% \def\file.../ beginning exact!

2

 3 %% A BibTeX bibliography style
 4 %%    conforming to the Chicago Manual's A style
 5 %%       but with B style author-date citations
 6 %%          by Matt Swift <swift@alum.mit.edu>
 7 %%
 8 %% \def\fileinfo{A Chicago Manual BibTeX style}
 9 %% \def\DoXPackageS {longtable}
10 %% \def\initelyHavECitationS {}
11 %% \def\fileversion{v1}
12 %% \def\filedate{2001/08/31}
13 %% \def\docdate{2001/08/31}
14 %%

15

4 Implementation

4.1 Define fields and variables

4.1.1 Global variables

output.state will have one of the four scalar values before.all, mid.sentence,

after.sentence, and after.block, telling us where we are in producing the entry.

Achicago uses only blocks, not sentences, to delimit units in the bibentry (some other bibstyles use both). The achicago package defines \newblock to add a very small and rubber amount of horizontal space, relaxing somewhat the \frenchspacing in effect. I have not removed the programming structures that distinguish sentences and blocks, but I have not concerned myself with them either; anyone wishing to revise this style to begin using them should check their work carefully.

16 INTEGERS { output.state before.all mid.sentence after.sentence after.block }

The entry variable principal-type will have one of the five scalar values

author.ptype, editor.ptype, organization.ptype, key.ptype, and emergency.ptype,

telling us what type of information is in the principal field. See also the discussion

of key.special.p.
17 INTEGERS { author.ptype editor.ptype organization.ptype key.ptype
18  emergency.ptype }

These are used during the forward and reverse passes to hold information about the previously-processed record.

19 INTEGERS { last.year.tag.num }
20 STRINGS  { last.leaditem last.label.and.year following.year.label.tag }

These variables are used locally within various functions or function groups, where their particular use is documented.

To do: document which if any are used in a nonlocal way

Warning: Suspect a conflict among these scratch variables if a strange problem occurs! I think u is the only one that will hold its value past an output.

21 INTEGERS { j nameptr namesleft numnames len }
22 STRINGS  { s t u v w }

4.1.2 Bibfields

The list of fields that will be examined in database records, if they exist. Fields in the bib database not on this list will be ignored. Not every entry type makes use of every field. Each entry type function below lists its required, optional, and ignored fields. Expected use of these fields is documented above in the user documentation.

23 ENTRY
24 { address
25  annotation
26  author
27  bookauthor
28  booktitle
29  chapter
30  edition
31  editor
32  faddress
33  flanguage
34  fpublisher
35  ftitle
36  fyear
37  howpublished
38  illustrator
39  institution
40  journal
41  key
42  month
43  note
44  number
45  oaddress
46  opublisher
47  organization
48  oyear
49  pages
50  publisher
51  school
52  series
53  seriesedition
54  title
55  translator
56  type
57  volume
58  year
59  yearcomp

60 }

Entry variables have a per-record value but are not initialized from a field.

The integer principal-type will contain one of the scalars author.ptype, editor.ptype, etc. (see their declarations above for further information). principal-type

is set in the function set.principal-field.principal-type.
61 { principal-type }

The strings label, year.label and year.label.tag are used to construct the citation (the running text at the location of the user’s \cite command). In most cases, label will contain a string derived from the principal-field (see

make.label).
62 { label

year.label will contain a string representing the date. It’s not necessarily the same as ‘’ year (see make.year.label). year.label.tag will either be empty or be a lower-case letter, as determined in the forward and reverse passes.

63  year.label year.label.tag

principal-field will contain one of the fields, depending on entry type and

what fields are nonempty (see set.principal-field.principal-type).

In most cases, leaditem will contain the principal-field properly formatted for apearance at the beginning of the bibliography entry (see make.leaditem). When it is identical to the previous entry’s leaditem, it will be wrapped in \SCduplicate by the forward and reverse passes.

64  principal-field
65  leaditem

66 }

4.2 Define macros

These are macro abbreviations required by convention in every bibstyle. The user can redefine them with @STRING declarations in the bib database.

67 MACRO {jan} {"January"}
68 MACRO {feb} {"February"}
69 MACRO {mar} {"March"}
70 MACRO {apr} {"April"}
71 MACRO {may} {"May"}
72 MACRO {jun} {"June"}
73 MACRO {jul} {"July"}
74 MACRO {aug} {"August"}
75 MACRO {sep} {"September"}
76 MACRO {oct} {"October"}
77 MACRO {nov} {"November"}
78 MACRO {dec} {"December"}

4.3 Debugging functions

debugval debugmsg
79 FUNCTION {debugval}
80 {  ": [" *
81 swap$ *    %$
82 "] in " * cite$ *    %$
83 top$      %$

84 }

 86 FUNCTION {debugmsg}
 87 {  " in " * cite$ * top$

88 }

4.4 Initialization functions

init.scalars This should be called before any processing begins.

 89 FUNCTION {init.scalars}
 90 { #0 'before.all :=
 91 #1 'mid.sentence :=
 92 #2 'after.sentence :=
 93 #3 'after.block :=

94

 95 #0 'author.ptype :=
 96 #1 'editor.ptype :=
 97 #2 'organization.ptype :=
 98 #3 'key.ptype :=
 99 #4 'emergency.ptype :=

100 }

4.5 Boolean logic

not and or
101 FUNCTION {not}
102 {   { #0

103 }

104  { #1

105 }

106 if$                 %$

107 }

108

109 FUNCTION {and}
110 {   'skip$            %$
111  { pop$            %$

112 #0 113 }

114 if$                 %$

115 }

116

117 FUNCTION {or}
118 {   { pop$            %$

119 #1 120 }

121  'skip$            %$
122 if$                 %$

123 }

4.6 Constant functions

comma period hyphen colon space delimiter.name-part delimiter.name delimiter.sub-sortkey tie etal unidentified These definitions are made so that later functions are easier to read and so that

adapting this bibstyle to another language will be simpler.

The reason these are functions not macros is so the user cannot change them.

124 FUNCTION {comma}
125 { ","

126 }

127

128 FUNCTION {period}
129 { "."

130 }

131

132 FUNCTION {hyphen}
133 { "-"

134 }

135

136 FUNCTION {colon}
137 { ":"

138 }

139

140 FUNCTION {space}
141 { " "

142 }

143

144 FUNCTION {delimiter.name-part}
145 { "  "

146 }

147

148 FUNCTION {delimiter.name}
149 { "   "

150 }

151

152 FUNCTION {delimiter.sub-sortkey}
153 { "    "

154 }

155

156 FUNCTION {tie}
157 { "~"

158 }

159

160 FUNCTION {etal}
161 { " et~al."

162 }

Below, we need to be able to set a string to a value which would never be generated from any string in a bibliography database. It’s safe to assume that this string is such a one.

163 FUNCTION {unidentified}
164 { "HiGHlY*Peco@lIEr"

165 }

4.7 Output functions

The output function automatically handles interunit punctuation.

output should be called with a string and a punctuation mark on the stack. The punctuation mark is what should precede the string in the case the string begins in the middle of a sentence. Strings which begin sentences are preceded by periods. Strings which begin a new block are preceded by periods and the \newblock command. (It makes no sense for a string to begin at the end of a sentence.)

Calls to output.internal leave the string on the stack until the next time output.internal is called, when it will get written. The function output.begin puts an empty string on the stack, and should be called before beginning any series of calls to output. output.end cleans up and should be called at the conclusion of a series of calls to output.

output.begin Prepare for a series of calls to output. Put an empty string on the stack and set

output.state to before.all.
166 FUNCTION{output.begin}
167 { ""
168 before.all 'output.state :=

169 }

output.end Conclude a series of calls to output. Simply calls write$.

170 FUNCTION {output.end}
171 { write$               %$

172 }

new.block Begin a new block.

173 FUNCTION {new.block}
174 { output.state before.all =
175  'skip$                  %$
176  { after.block 'output.state :=

177 }

178 if$                       %$

179 }

new.sentence Begin a new sentence. FIX: hmm, wouldn’t we want to put a TeX command here for intersentence space now? Why do we mess with the usual spacefactor for periods anyway? When in bibliogrphies are the usual TeX rules not going to be what we want, and give too much space after a period? answer: when a lowercase letter followed by period, in e.g. sec. 2 and so on. often bibstyle will put tie there, but sometimes not I think, so need to have that be not an intersentence space. well, if i start using sentences instead of blocks, I think perhaps doing the same hskip after a sentence with a new command \newsentence perhaps as in \newblock would be the right thing to do.

180 FUNCTION {new.sentence}
181 { output.state before.all = output.state after.block = or
182  'skip$              %$
183  { after.sentence 'output.state :=

184 }

185 if$              %$

186 }

output.internal \stackpunct string/entry(nonnull)null

187 FUNCTION {output.internal}
188 { 's :=     % s := string/entry(nonnull)
189 space * 't :=     % t := punctuation + space
190 output.state mid.sentence =

191  { t * write$               %$

192 }

193  { output.state after.block =
194  { add.period$ write$
195  newline$              %$
196  "\newblock " write$              %$

197 }

198  { output.state before.all =
199  'write$              %$

Final case is output.state = ‘after.sentence’, which means we just add a period, but no newline or \newblock command.

200  { add.period$ space * write$

201 }

202  if$              %$

203 }

204  if$              %$

We are now mid-sentence, whether we were before or not.

205  mid.sentence 'output.state :=

206 }

207 if$              %$

The actual string given to ‘output.internal’ is left on the stack here for next time we call ‘output.internal’ or ‘output.end’, whichever comes next.

208 s 209 }

output \stackstring/entry puncteither null OR punct string/entry(nonnull)

Check if string/entry is null. If it’s null do nothing. If it’s nonempty call output.internal on it with the given punctuation.

210 FUNCTION {output}
211 { swap$ duplicate$ empty$
212  { pop$
213  pop$              %$

214 }

215  'output.internal
216 if$              %$

217 }

output.bibitem The optional argument to \bibitem will be made the definition of \b@foo where foo is the cite$ key. \b@foo will be executed by all the various citing commands with a definition of \SCcite that is appropriate to each citing command.

The thebibliography environment actually ignores the optional argument to \bibitem. The environment sets the formatting so that the first part of the main entry hangs out to the left and looks like a label.

The unexpandable protection is necessary, instead of a simple protection, in order to work with newclude’s tag option.

I guess we use write$ for each part instead of catenating them and using a single write$ in case the catenation gets longer than BibTEX can handle. I don’t know, this is how I found it. It seems like catenation and a single call to write$ would be faster. FIX.

218 FUNCTION {output.bibitem}
219 { newline$                 %$

220 "\bibitem[\UnexpandableProtect\SCcite{" write$          %$
221 label write$           %$
222 "}{" write$                  %$
223 year.label year.label.tag * write$        %$
224 "}]{" write$             %$
225 cite$ write$
226 "}" write$               %$
227 newline$            %$

228 }

4.8 Simple auxiliary functions

field.or.null \stackfield literalfield contents or null string if the field was missing

229 FUNCTION {field.or.null}
230 { duplicate$ empty$
231  { pop$             %$

232 "" 233 }

234  'skip$             %$
235 if$                  %$

236 }

italicize \stackstring or entrystring

Italicize the top string on the stack and provide italic corrections. Null goes to null. We want to explicitly italicize rather than use \emph, since semantic emphasis might be indicated in some other way in the document, e.g., slanted type.

237 FUNCTION {italicize}
238 { duplicate$ empty$
239  { pop$              %$

240 "" 241 }

242  { "\textitswitch{" swap$ * "}" *       %$

243 }

244 if$                       %$

245 }

parenthesize \stackstring-or-entry:S”(” + S + ”)”

246 FUNCTION {parenthesize}
247 { duplicate$ empty$
248  { pop$            %$

249 "" 250 }

251  { "(" swap$ * ")" *         %$

252 }

253 if$                           %$

254 }

format.clip \stackstringstring FIX: the difference between entry.max$ and global.max$: max for an entry string, max for a global string.

format.clip \stackstring-Sstring-S-possibly-shortened Clips a string down to the smaller of entry.max$ and global.max$.

255 FUNCTION {format.clip}
256 { #1 entry.max$ substring$
257 #1 global.max$ substring$

258 }

format.sortify.clip purify$, downcase, and clip a string on the stack. \stack⟨string⟩⟨string⟩

259 FUNCTION {format.sortify.clip}
260 { % duplicate$ "sinful" debugval %$
261 purify$                   %$
262 %  duplicate$ "  pure" debugval %$
263 "l" change.case$                   %$
264 format.clip
265 %  duplicate$ "humble" debugval %$

266 }

check.empty \stackS warningS Issues warning if the string is empty.

267 FUNCTION {check.empty}
268 { 't :=
269 duplicate$ empty$
270  { "empty " t * " in " * cite$ * warning$

271 }

272  'skip$                              %$
273 if$                              %$

274 }

multipage.p  \stacknothing#0 or #1

Set t to ‘’ pages. Look at t character by character for a hyphen, comma, or plus. If we find one, return true immediately; else return false.

275 FUNCTION {multipage.p}
276 { pages 't :=
277 #0 'j :=
278 % WHILE
279 {   j not
280  t empty$ not            %$
281  and

282 }

283  { t #1 #1 substring$                %$
284  duplicate$ hyphen =                  %$
285  swap$ duplicate$ comma =
286  swap$ "+" =                       %$

287 or 288 or

289  { #1 'j :=

290 }

291  { t #2 global.max$ substring$ 't :=

292 }

293  if$

294 }

295 while$

296 j 297 }

format.n.dashify  \stackstringstring

Make single-dashes into double-dashes.

298 FUNCTION {format.n.dashify}
299 { 't :=

300 ""

301 % WHILE
302 { t empty$ not }                %$
303  { t #1 #1 substring$ hyphen =                %$
304  { t #1 #2 substring$ "--" = not                %$
305  { "--" *
306  t #2 global.max$ substring$ 't :=

307 }

308  { % WHILE
309  { t #1 #1 substring$ hyphen = }                %$
310 { hyphen *
311 t #2 global.max$ substring$ 't :=

312 }

313  while$                %$

314 }

315 if$                %$

316 }

317  { t #1 #1 substring$ *                %$
318  t #2 global.max$ substring$ 't :=

319 }

320  if$                %$

321 }

322 while$                                  %$

323 }

tie.or.space.connect  \stackstring stringstring

If the second string has less than 5 characters, the two are tied together, otherwise they are catenated with a space between them.

324 FUNCTION {tie.or.space.connect}
325 { duplicate$ text.length$ #5 <
326  { tie

327 }

328  { space

329 }

330 if$
331 swap$ * *

332 }

either.or.check \stackmsg fieldnull If field is not empty, give a warning appropriate for saying this field and an already-existing field can’t be both used in the same entry.

333 FUNCTION {either.or.check}
334 { empty$
335  'pop$
336  { "can't use both " swap$ * " fields in " * cite$ * warning$

337 }

338 if$

339 }

4.9 Queue functions START FIXING DOX HERE

Queue functions expect a string on the stack. If it is empty, and the field for which the function is named is noempty, the string will be replaced with that field. (e.g., author.q will put ‘’ author on the stack if ‘’ author is nonempty and the author.q was passed an empty string). A chain of queue functions must begin with something on the stack, adn will leave the first nonempty field named in the chain on the stack.

Queue functions also increment an integer variable, ’j, whenever they are passed an empty string. At the end of the queue, ’j tells us which function in the queue contributed the string that results at the end of the queue, starting with zero for the first queue function after start.a.q. In order for this to work, queue functions must always be called in the same order: ‘’ author (’#0’) ‘’ editor (‘#1’), ‘’ organization (‘#2’), and ‘’ key (‘#3’). For entry types where a certain field should not appear in the queue, use ‘blank.q’ in its place in the order.

The value of j is consulted when setting leaditem, but not in other cases where queue functions are used

start.a.q must begin a queue. It initializes j and passes an empty string to the first queue function.

start.a.q
340 FUNCTION {start.a.q}
341 { #-1 'j :=

342 "" 343 }

inc.j
key.q
organization.q
editor.q
author.q
blank.q

344 FUNCTION {inc.j}
345 { j #1 + 'j :=

346 }

347

348 FUNCTION {key.q}
349 { duplicate$ empty$
350  { pop$           %$
351  key field.or.null
352  inc.j

353 }

354  'skip$             %$
355 if$                  %$

356 }

357

358 FUNCTION {organization.q}
359 { duplicate$ empty$
360  { pop$                  %$
361  organization field.or.null
362  inc.j

363 }

364  'skip$                    %$
365 if$                       %$

366 }

367

368 FUNCTION {editor.q}
369 { duplicate$ empty$

370  { pop$                       %$
371  editor field.or.null
372  inc.j

373 }

374  'skip$                       %$
375 if$                       %$

376 }

377

378 FUNCTION {author.q}
379 { duplicate$ empty$
380  { pop$                       %$
381  author field.or.null
382  inc.j

383 }

384  'skip$                       %$
385 if$                       %$

386 }

387

388 FUNCTION {blank.q}
389 { duplicate$ empty$
390  'inc.j
391  'skip$                       %$
392 if$                       %$

393 }

check.q \stackstring-S warning-stringstring-S

This should be the final queue command, though it is not a true queue command, since an extra warning string should be on the top of the stack above the string that has passed along the queue. The warning string should be an English disjunctive list of the queue commands that have come before it, so the user can tell which fields are missing. For example “author, organization, or key” if the queue has been author.q ’blank.q’ ‘organization.q’ key.q.

If the queue string is empty, check.q issues the warning and puts on the stack an emergency string of the L^{A}TEX key (i.e., cite$), which should allow BibTEX to proceed.

394 FUNCTION {check.q}
395 { 't :=
396 duplicate$ empty$
397  { pop$                              %$
398  cite$ format.clip 'u :=           %$
399  "Need " t * " in " * cite$ * "; using:  " * u * warning$

400 u

401  inc.j

402 }

403  'skip$                              %$
404 if$                              %$

405 }

4.10 Format functions dox

The format functions modify a string on the stack. They make no warnings, produce no output, and place no spaces or punctuation before or after their string, just modify it somehow.

4.11 Format names functions

format.names.for.label Format string on stack it to produce something appropriate for label. Comma usage examples in the Chicago Manual: “A et al.” and “A, B, et al.” The Chicago Manual recommends for labels: use up to 3 names, and above

3 use 1st + etal. When this would lead to ambiguity, use one of 3 solns: use all names, use as many as needed to disambiguate, or include a (short) title to disambiguate: (Zipursky et al., Brief Title, 1988).

The Chicago Manual does not appear to distinguish between “et al.” and “and others”. It’s possible to change the logic here to distinguish between the cases where the bibstyle is truncating the list of authors, and the case where “and others” appears in the bib file. But I’m not doing that now.

Labels:

FIX: this stuff not implemented yet 1 author Foo
2 authors  Foo and Bar
3 authors  Foo, Bar, and Baz
4+ authors  Foo et al.
1+ author “and others”  Foo et al.

This function is probably more general than it needs to be, but it’s a nice model.

406 FUNCTION {format.names.for.label}
407 { 's :=
408 s num.names$ 'numnames :=                                   %$
409 s #1 "{vv~}{ll}" format.name$                                 %$
410  s numnames "{ff}{vv}{ll}{jj}" format.name$ "others" =    %$
411  numnames #3 >

412 or

413  % more than 3 names, or final name is "others"
414  { etal *

415 }

416  { numnames #1 - 'namesleft :=
417  #2 'nameptr :=
418 % WHILE
419 { namesleft #0 > }
420 { nameptr numnames =
421  % we're on the final name (and it's not "others")
422  { numnames #2 >
423 { comma *

424 }

425 'skip$        %$
426  if$        %$
427  " and " * s nameptr "{vv~}{ll}" format.name$ *     %$

428 }

429  % we're not on the final name (in this case: 2d of 3 names)
430  { comma * space *
431  s nameptr "{vv~}{ll}" format.name$ *        %$

432 }

433 if$          %$
434 nameptr #1 + 'nameptr :=
435 namesleft #1 - 'namesleft :=

436 }

437  while$             %$

439 if$                    %$

440 }

format.names.lastfirst Pop a string containing a name list. Push a string with the first author last-name-first and subsequent authors first-name-first.

441 FUNCTION {format.names.lastfirst}
442 { 's :=
443 s num.names$ 'numnames :=    % numnames = num.name$(s);
444 numnames 'namesleft :=
445 #1 'nameptr :=               % nameptr = 1;
446  % WHILE
447  { namesleft #0 > }
448  { nameptr #1 =
449  { s nameptr "{vv~}{ll}{, jj}{, ff}" format.name$                %$

450 }

451  { comma *
452  s nameptr "{ff~}{vv~}{ll}{, jj}" format.name$ 't :=           %$
453  nameptr numnames =
454  % final name
455  { t "others" =
456  { etal *

457 }

458  { " and " * t *

459 }

460  if$                  %$

461 }

462  % names both precede and follow this name
463  { space * t *

464 }

465  if$                  %$

466 }

467  if$                  %$
468  nameptr #1 + 'nameptr :=          % nameptr += 1;
469  namesleft #1 - 'namesleft :=      % namesleft =- 1;

470 }

471 while$                     %$

472 }

format.names.firstfirst Same as format.names.lastfirst but all names are first name first.
473 FUNCTION {format.names.firstfirst}
474 { 's :=
475 s num.names$ 'numnames :=    % numnames = num.name$(s);
476 numnames 'namesleft :=
477 #1 'nameptr :=               % nameptr = 1;
478  % WHILE
479  { namesleft #0 > }
480  { s nameptr "{ff~}{vv~}{ll}{, jj}" format.name$ 't :=                   %$
481  nameptr #1 =

482 't

483  { nameptr numnames =
484  % final name
485  { numnames #2 >
486  { comma *

488  'skip$                  %$
489  if$                  %$
490  t "others" =
491  { etal *

492 }

493  { " and " * t *

494 }

495  if$                  %$

496 }

497  % names both precede and follow this name
498  { comma * space * t *

499 }

500  if$                  %$

501 }

502  if$                  %$
503  nameptr #1 + 'nameptr :=          % nameptr += 1;
504  namesleft #1 - 'namesleft :=      % namesleft =- 1;

505 }

506 while$                   %$

507 }

4.12 Make functions

Make functions expect nothing on the stack, and leave one string there. Any side effects of setting global variables are noted in the comments. If a nonempty string cannot be made due to lack of the right fields, leave an empty one on the stack.

make.annotation Annotations can be omitted by nullifying the SCannotation environment.

508 FUNCTION {make.annotation}
509 { annotation empty$       %$
510  { ""

511 }

512  { "\begin{SCannotation}"
513  annotation *
514  "\end{SCannotation}" *

515 }

516 if$           %$

517 }

make.edition We use some logic here because we want to use the ‘’ seriesedition for the ‘’ edition in certain cases.

518 FUNCTION {make.edition}
519 { edition empty$                  %$
520  { seriesedition empty$ not series field.or.null title field.or.null = and             %$
521  { seriesedition

522 }

523  { ""

524 }

525  if$                                    %$

526 }

527  { edition

528 }

529 if$                                        %$

530 }

make.year.month make.year.nomonth Make a year string. We don’t we use year.label because that string is sterilized to make it appropriate for use in running text in citations (and for sorting I think); in the bibliography entry, we can be a little more lax, accepting special characters, and text that is not simply four numerals.

531 FUNCTION {make.year.month}
532 { year empty$                  %$
533  { ""

534 }

535  { year year.label.tag *
536 month empty$                  %$
537 'skip$                  %$
538 { comma * space * month *

539 }

540 if$                  %$

541 }

542  if$                                      %$

543 }

544 FUNCTION {make.year.nomonth}
545 { year empty$                                %$
546  { ""

547 }

548  { year year.label.tag *

549 }

550 if$                                        %$

551 }

make.year.or.oyear.month Push a string with year information. If ‘’ oyear is empty, call make.year.month. Warn in the almost-certainly-erroneous case of an ‘’ oyear without a ‘’ year.

552 FUNCTION {make.year.or.oyear.month}
553 { oyear empty$                  %$
554  { make.year.month

555 }

556  { year empty$                  %$
557  { "oyear without year in " cite$ * warning$

558 }

559  'skip$                  %$
560  if$                  %$
561  oyear year.label.tag *
562  month empty$                  %$
563  'skip$                  %$
564 { comma * space * month *

565 }

566  if$                  %$

567 }

568 if$                    %$

569 }

make.leaditem.with.tags Push leaditem and if it’s an editor, include a nice taggy.

570 FUNCTION {make.leaditem.with.tags}
571 { leaditem

572 principal-type editor.ptype =
573  { editor num.names$ #1 >         %$
574 { ",~eds." *

575 }

576 { ",~ed." *

577 }

578  if$                  %$

579 }

580  'skip$                 %$
581 if$                 %$

582 }

make.editors.secondary Give editors’ names when not for the leaditem.

583 FUNCTION {make.editors.secondary}
584 { editor empty$                  %$
585  { ""

586 }

587  { output.state after.sentence =
588  output.state after.block = or
589  { "E"

590 }

591  { "e"

592 }

593  if$                  %$
594  "dited by " * editor format.names.firstfirst *

595 }

596 if$                                        %$

597 }

make.bookauthors.secondary This is for making authors’ names that follow a title. When we use this, we are only going to be interested in the ‘’ bookauthor.

598 FUNCTION {make.bookauthors.secondary}
599 { bookauthor empty$                  %$
600  { ""

601 }

602  {   output.state after.sentence =
603  output.state after.block =

604 or

605  { "B"

606 }

607  { "b"

608 }

609  if$                  %$
610  "y " * bookauthor format.names.firstfirst *

611 }

612 if$                  %$

613 }

make.translators make.illustrators Translators and illustrators are always secondary, until I write some extensions.

614 FUNCTION {make.translators}
615 { translator empty$                  %$
616  { ""

617 }

618  {   output.state after.sentence =

619  output.state after.block =

620 or

621  { "T"

622 }

623  { "t"

624 }

625  if$                  %$
626  "ranslated " *
627  flanguage empty$                  %$
628  'skip$                  %$
629  { "from the " * flanguage * space *

630 }

631  if$                  %$
632  "by " * translator format.names.firstfirst *

633 }

634 if$                  %$

635 }

636

637 FUNCTION {make.illustrators}
638 { illustrator empty$                  %$
639  { ""

640 }

641  {   output.state after.sentence =
642  output.state after.block =

643 or

644  { "I"

645 }

646  { "i"

647 }

648  if$                  %$
649  "llustrated by " * illustrator format.names.firstfirst *

650 }

651 if$                  %$

652 }

make.booktitle.ital make.title.ital make.as.ftitle.ital FIX: multilingual
653 FUNCTION {make.title.ital}
654 { title field.or.null italicize

655 }

656

657 FUNCTION {make.booktitle.ital}
658 { booktitle field.or.null italicize

659 }

660

661 FUNCTION {make.as.ftitle.ital}
662 { ftitle empty$                  %$
663  { ""

664 }

665  { "as " ftitle italicize *

666 }

667 if$                  %$

668 }

make.title.inquotes Put the title in quotes on the stack. \Wrapquotes is going to properly handle punctuation that follows the closing quotation marks by sucking it inside the quotes. FIX: make generic function format.wrapquotes? like format.emphasize? FIX: make all those function that transform a string on the stack into “format” functions.

669 FUNCTION {make.title.inquotes}
670 { title empty$               %$
671  { ""

672 } 673 {

674  "\Wrapquotes{" title * "}" *

675 }

676 if$                                  %$

677 }

678 FUNCTION {make.pages}
679 { pages empty$               %$
680  { ""

681 }

682  { pages format.n.dashify

683 }

684 if$                                     %$

685 }

make.vol.series.num.month.pages For articles, the number of cases to handle these fields is very large: 32 to do it properly! These cases could perhaps be collapsed, but it is going to be much easier to read here and to maintain if we do it brutishly by elaborating each case. The only exception I’ve made is when two adjacent leaves are very similar. The way I’ve laid out this funciton here, adjacent leaves are going to be the pairs of cases with and without a nonempty ‘’ pages.

There are 5 relevant fields, each of which can be empty or not (2^{5} = 32). In the comments below, I mark each case with a series of 1 to 5 letters represent the cases where that field is nonempty: V for ‘’ volume, S for ‘’ series, N for ‘’ number, M for ‘’ month, and P for ‘’ pages. Thus, the case marked with the comment SMP is the one where the ‘’ series, ‘’ month, and ‘’ pages fields are nonempty, and ‘’ volume and ‘’ number are empty.

FIX: when I use make.pages in this function, it’s slightly inefficient, since make.pages does a check for empty pages which is unnecessary here.

FIX: this does not yet handle entries with nonempty ‘’ series.

FIX: go through and combine adjacent leaves.

We could really put the punctuation that should precede this unit right at the beginning of the string we’re making here, but let’s conform better to the conventions and put this in the scratch variable ’u instead. I don’t think this function will ever get called at any time other than mid-sentence, after the journal title, but who knows. So a call to this function should look like:

   make.vol.series.num.month.pages u output

Remember that when output.internal writes the punctuation, it adds a space following, so it is unnecessary to begin the string this function returns with a space. (Occasionally below, we want (and produce) a colon followed by no space; we can do this because we are directly creating this string, not using the output routine.)

686 FUNCTION {make.vol.series.num.month.pages}

687 { volume empty$                   %$
688  { series empty$                   %$
689  { number empty$                   %$
690  { month empty$                   %$
691 { pages empty$                   %$
692  % none of VSNMP
693  { space 'u :=

694 "" 695 } 696 % P

697  { comma 'u :=
698  multipage.p
699  { "pp." pages format.n.dashify

700 }

701  { "p." pages

702 }

703  if$                   %$
704  tie.or.space.connect

705 }

706 if$                   %$

707 }

708 { pages empty$                   %$

709 % M

710  { comma 'u :=
711  month

712 }

713  % MP
714  { comma 'u :=
715  month comma * space * make.pages *

716 }

717 if$                   %$

718 }

719  if$                   %$

720 }

721  { month empty$                   %$
722 { pages empty$                   %$

723 % N

724  { comma 'u :=
725  "no." number tie.or.space.connect

726 }

727  % NP
728  { comma 'u :=
729  "no." number colon * make.pages * tie.or.space.connect

730 }

731 if$                   %$

732 }

733 { pages empty$                   %$
734  % NM
735  { comma 'u :=
736  "no." number tie.or.space.connect
737  space * month parenthesize *

738 }

739  % NMP
740  { comma 'u :=

741  "no." number tie.or.space.connect
742  space * month parenthesize *
743  colon * space * make.pages *

744 }

745 if$                   %$

746 }

747  if$                   %$

748 }

749 if$                   %$

750 }

751  { number empty$                   %$
752  { month empty$                   %$
753 { pages empty$                   %$

754 % S 755 { 756 }

757  % SP

758 { 759 }

760 if$                   %$

761 }

762 { pages empty$                   %$
763  % SM

764 { 765 }

766  % SMP

767 { 768 }

769 if$                   %$

770 }

771  if$                   %$

772 }

773  { month empty$                   %$
774 { pages empty$                   %$
775  % SN

776 { 777 }

778  % SNP

779 { 780 }

781 if$                   %$

782 }

783 { pages empty$                   %$
784  % SNP

785 { 786 }

787  % SNMP

788 { 789 }

790 if$                   %$

791 }

792  if$                   %$

793 }

794 if$                   %$

796  if$                   %$

797 }

798  { series empty$                   %$
799  { number empty$                   %$
800  { month empty$                   %$
801 { pages empty$                   %$

802 % V

803  { comma 'u :=
804  "vol." volume tie.or.space.connect

805 }

806  % VP
807  { space 'u :=
808  volume colon * make.pages *

809 }

810 if$                   %$

811 }

812 { pages empty$                   %$
813  % VM
814  { comma 'u :=
815  "vol." volume tie.or.space.connect
816  space * month parenthesize *

817 }

818  % VMP
819  { space 'u :=
820  volume space *
821  month parenthesize * colon * space * make.pages *

822 }

823 if$                   %$

824 }

825  if$                   %$

826 }

827  { month empty$                   %$
828 { pages empty$                   %$
829  % VN
830  { space 'u :=
831  volume ", no." number tie.or.space.connect *

832 }

833  % VNP
834  { space 'u :=
835  volume space *
836  number parenthesize * colon * space * make.pages *

837 }

838 if$                   %$

839 }

840 { pages empty$                   %$
841  % VNM
842  { space 'u :=
843  volume ", no." number tie.or.space.connect *
844  space * month parenthesize *

845 }

846  % VNMP
847  { space 'u :=
848  volume space * number parenthesize * colon * space *

849  make.pages * space * month parenthesize *

850 }

851 if$                   %$

852 }

853  if$                   %$

854 }

855 if$                   %$

856 }

857  { number empty$                   %$
858  { month empty$                   %$
859 { pages empty$                   %$
860  % VS

861 { 862 }

863  % VSP

864 { 865 }

866 if$                   %$

867 }

868 { pages empty$                   %$
869  % VSM

870 { 871 }

872  % VSMP

873 { 874 }

875 if$                   %$

876 }

877  if$                   %$

878 }

879  { month empty$                   %$
880 { pages empty$                   %$
881  % VSN

882 { 883 }

884  % VSNP

885 { 886 }

887 if$                   %$

888 }

889 { pages empty$                   %$
890  % VSNM

891 { 892 }

893  % VSNMP

894 { 895 }

896 if$                   %$

897 }

898  if$                   %$

899 }

900 if$                   %$

901 }

902  if$                   %$

904 if$                                            %$

905 }

make.crossref
906 FUNCTION {make.crossref}
907 { "\citeNP{" crossref * "}" *

908 }

output.series.number Outputs the number with appropriate capitalization. Warns if we also have nonempty ‘’ seriesedition.

909 FUNCTION {output.series.number}
910 { series empty$                   %$
911  { number empty$                   %$
912  { ""

913 }

914  { "Can't use `number' without `series' in " cite$ * warning$

915 "" 916 }

917  if$                   %$

918 }

919  { number empty$                   %$
920 { series
921  seriesedition empty$                   %$
922  'skip$                   %$
923  { comma output
924  new.block
925  seriesedition

926 }

927  if$                   %$

928 }

929 { series space *
930  "no." *
931 number tie.or.space.connect
932 "number and seriesedition" seriesedition either.or.check

933 }

934  if$                   %$

935 }

936 if$                                         %$
937 comma output

938 }

output.volume.series The ‘’ series is not required. FIX: why do we output stuff here? can’t we use a make function like elsewhere?

939 FUNCTION {output.volume.series}
940 { volume empty$                   %$
941  { ""

942 }

943  { "Volume" volume tie.or.space.connect
944  series empty$                   %$
945  'skip$                   %$
946  { " of " * series italicize *
947  seriesedition empty$                   %$
948  'skip$                   %$

949  { comma output
950  new.block
951  seriesedition

952 }

953  if$                   %$

954 }

955  if$                   %$
956  "volume and number" number either.or.check

957 }

958 if$                   %$
959 comma output

960 }

make.chapter.of To say for example “Part 3” instead of “Chapter 3”, put “Part” into the ‘’ type field, and the number and the preposition into the ‘’ chapter field, e.g., “3 of”. A tie is automatically put in between the ‘’ type field and the ‘’ chapter field in this case. We can’t make a decision based on the length of the ‘’ chapter field when it’s going to contain the preposition.

OK, putting the preposition in the chapter field doesn’t work out when you do crossreferences. So let’s try “of” as a general preposition.

961 FUNCTION {make.chapter.of}
962 { chapter empty$        %$
963  { ""

964 }

965  { type empty$        %$
966  { "Chapter"

967 }

968  { type "t" change.case$         %$

969 }

970  if$        %$
971  chapter tie.or.space.connect
972  " of " *

973 }

974 if$                   %$

975 }

make.chapter
976 FUNCTION {make.chapter}
977 { chapter empty$                   %$
978  { ""

979 }

980  { type empty$                   %$
981  { "Chapter"

982 }

983  { type "t" change.case$                   %$

984 }

985  if$                   %$
986  chapter tie.or.space.connect

987 }

988 if$                   %$

989 }

make.number.tr

990 FUNCTION {make.number.tr}
991 { type empty$                   %$
992  { "Technical Report"

993 }

994  { type "t" change.case$                    %$

995 }

996 if$                   %$
997 number empty$                   %$
998  'skip$                   %$
999  { number tie.or.space.connect

1000 }

1001 if$                   %$

1002 }

4.13 Label and principal field functions

calc.labels set.principal-field.principal-type Set label and year.label from the original database. Set leaditem. For refer-

ence from above: ‘’ author (’#0’) ‘’ editor (‘#1’), ‘’ organization (‘#2’), and ‘’ key (‘#3’).

FIX: NOTE: the entry-type functions (and specifically its subroutines like make.authors) are going to assume that bibentry contains the right field as calculated by

format the string on the stack based on the GLOBAL (not entry) variable j (= bibentry.type)

confusion here over terminology. bibentry and leaditem are the same thing, but it is also the basis for calculating the label. let’s call the most important field for each type the PRINCIPAL field. this will be put into the lead position in the bibentry and also be massaged into the label. FIX: new idea, start using it leaves a string on the stack based on TYPE; the string is the first non-empty relevant field for the type, or an emergency substitution as supplied by check.q. FIX: end.the.q? would emergency be a 5th type that the format.principal-field functions should recognize?

Yes, check.q will put full cite-string on as emergency and set principal-type accordingly.

FIX: ensure proper warnings if we are not getting something good in principalfield.

1003 FUNCTION {set.principal-field.principal-type}
1004 {   type$ "book" =               %$
1005  type$ "inbook" =               %$

1006 or

1007  { start.a.q
1008  author.q
1009  editor.q
1010  blank.q
1011  key.q
1012  "author, editor, or key for entry type " type$ * check.q          %$

1013 }

1014  { type$ "proceedings" =                   %$
1015  { start.a.q
1016  blank.q
1017  editor.q
1018  organization.q

1019  key.q
1020  "editor, organization, or key for entry type " type$ * check.q    %$

1021 }

1022  { type$ "manual" =                   %$
1023  { start.a.q
1024  author.q
1025  blank.q
1026  organization.q
1027  key.q
1028  "author, organization, or key for entry type " type$ * check.q   %$

1029 }

1030  % for all other entry types
1031  { start.a.q
1032  author.q
1033  blank.q
1034  blank.q
1035  key.q
1036  "author or key for entry type " type$ * check.q       %$

1037 }

1038  if$                   %$

1039 }

1040  if$                   %$

1041 }

1042 if$                   %$
1043 'principal-field :=
1044 j 'principal-type :=

1045 }

1046

1047 % TRUE if we've got \evar{principal-type} of \gvar{organization.ptype} and
1048 % nonempty \field{key} and entry type of \entry{manual} or \entry{proceedings}.
1049 FUNCTION {key.special.p}
1050 {       type$ "manual" =             %$
1051 type$ "proceedings" =        %$

1052 or

1053  principal-type organization.ptype =

1054 and

1055  key empty$ not                 %$

1056 and 1057 }

1058

1059 FUNCTION {make.leaditem}
1060 { key.special.p
1061  % is \field{key} all-caps?
1062  { key duplicate$ "u" change.case$ =
1063  % if so, key is assumed to be an acronym FIX; should I warn what we're doing?
1064  { key " (" * organization * ")" *

1065 }

1066  % if not, key is assumed to be an initial prefix (abbreviation) of organization
1067  % for use only in the label
1068  { organization

1069 }

1070  if$                   %$

1071 }

1072  { principal-field

1073 principal-type author.ptype =
1074 principal-type editor.ptype =

1075 or

1076 { format.names.lastfirst

1077 }

1078 % organization, key, or emergency ptype
1079 'skip$               %$
1080  if$                   %$

1081 }

1082 if$              %$

1083 }

1084

1085 FUNCTION {make.label.from.principal-field}
1086 { principal-field
1087  principal-type author.ptype =
1088  principal-type editor.ptype =

1089 or

1090  { format.names.for.label

1091 }

1092  % organization, key, or emergency
1093  'skip$               %$
1094 if$                   %$

1095 }

1096

1097 FUNCTION {make.label}
1098 { key.special.p
1099  { key

1100 }

1101  { make.label.from.principal-field

1102 }

1103 if$               %$

1104 }

format.names.for.sort \stack⟨list of names⟩⟨list of names suitable for sorting⟩ FIX: add dox here about how bibtex sorts, and the spaces strategy.

1 between tokens of a name part (word-by-word method; only af- ter first-name’s initials in letter- by-letter) 2 between name- parts in a name (but consider ‘von’ and ‘last’ one name-part) 3 between names in a name list 4 between other sort levels, such as names and years We don’t bother to format.sortify.clip, just do that once before assignment to the sort.key.

Chicago Manual specifies all names should be given in the bibliography. FIX ref. So the Chicago Manual doesn’t tell us how to handle the conventional “and others” that might be in the database. We handle it by proceeding on the assumption that “and others” always means the total names are more than 3, so we want to alphabetize these entries after all those with 1, 2, or 3 names. We do this by turning “and others” into a pseudo-name that will sort later than any real name (i.e., we insert a lot of z’s).

The commented-out format is doing word-by-word alphabetization as defined §17.97 with the exception of hyphens. purify$ turns hyphens into spaces, so hyphens are word boundaries, unlike slashes and apostrophes, which are suppressed. The Chicago Manual says hyphens should be supressed like slashes and apostrophes, so we need to remove them with format.dehyphen before we purify$.

The Chicago Manual prefers letter-by-letter alphabetization, which is what the active code below does. We suppress spaces between tokens of a name part. We need to remove hyphens as explained above.

NOTE: We could therefore change our delimiters all down by one, since single space delimters are no longer used, but I think there is no harm in starting with a 2-space delimiter, and this lets us change back and forth easily.

FIX: §17.100 names with initials in place of given names should precede those names which are spelled out. “J. Robert” before “James N.” and “K. T.” before “Keven S.”. This is not consistent with the letter-by-letter method (it is with word-by-word). So, we want to insert a space after first-name tokens of length 1. How do we do that!?

FIX: do we need special handling for names with Saint, Mac, Mc, and M’ §17.107–109?

1105

1106 % FIX: NOTE: \builtin{empty} returns TRUE if string contains no non-whitespace chars
1107 % FIX: other whitespace to check for?

1108

1109 % pop string, push its length not counting special chars as one char
1110 FUNCTION {length}
1111 { 'u :=
1112 %  u "taking length of" debugval
1113 #0 'len :=
1114 % WHILE
1115 { u "" = not }
1116  { len #1 + 'len :=
1117  u #2 global.max$ substring$ 'u :=

1118 }

1119 while$         %$

1120 len

1121 %  len int.to.str$ "      length is" debugval    %$

1122 }

1123 1124

1125 % FIX: a special+ char begins with a |{\| and continues through the next |}|.
1126 % \BibTeX{}'s ``special character'' is the same but only at brace level~0.
1127 % When we're stepping through a string, and we don't keep track of the brace
1128 % level, then those functions that recognize special chars are not going to
1129 % know the brace level, they will see every special+ char as a special char.

1130

1131 % pop string, push its first char counting special chars as one char
1132 % treat braces normally unless they are the start of a special char

1133 %

1134 % don't use \function{and} here for efficiency
1135 FUNCTION {head}
1136 { 'w :=
1137 w #1 #1 substring$      %$
1138 duplicate$ "{" =       %$
1139  { w #2 #1 substring$ "\" =      %$
1140  { pop$       %$
1141  w #1 text.prefix$     %$

1142 }

1143  'skip$   %$
1144  if$       %$

1145 }

1146  'skip$      %$
1147 if$        %$

1148 }

1149

1150 % FIX: using `and' and `or' requires calculation of both where nested branches
1151 % would not; see where this will help.

1152

1153 % pop string, push string without its first char, counting special chars as one char
1154 FUNCTION {behead}
1155 { 'u :=
1156 %  u "beheading before" debugval

1157 u

1158  u head length #1 +
1159  global.max$            %$
1160 substring$      %$
1161 %  duplicate$ "beheading  after" debugval    %$

1162 }

1163

1164 FUNCTION {behead.t}
1165 { t behead 't :=

1166 }

1167

1168 FUNCTION {format.dehyphen}
1169 { 't :=
1170 %  t "dehyphen before" debugval

1171 ""

1172 % WHILE
1173 { t "" = not }
1174  { t head
1175  duplicate$ hyphen =         %$
1176  { pop$                    %$

1177 } 1178 { * 1179 }

1180  if$                %$
1181  behead.t

1182 }

1183 while$                                  %$
1184 %  duplicate$ "dehyphen after" debugval    %$

1186

1187 % \builtin{text.length}: replace string on stack with its length, special chars count
1188 % as 1 char, don't count braces.

1189

1190 % \builtin{text.prefix}: like \builtin{substring} but count special chars as 1 char, don't
1191 % count braces, and supply a missing closing brace for a special-char.  also,
1192 % have to always start at \#1.

1193

1194 % FIX: does \manual intend to skip over en-dash and em-dash?  these could occur
1195 % in titles.  For the moment I will remove them too, but not in special chars.
1196 % look at \function{format.n.dashify} to be able to distinguish hyphens from
1197 % the dashes.

1198

1199 FUNCTION {format.dehyphen.dewhitespace}
1200 { 't :=
1201 %  t "dehyphen.dewhitespace before" debugval

1202 ""

1203 % WHILE
1204 { t "" = not }
1205  { t head
1206  duplicate$                  %$
1207  duplicate$ hyphen =       %$
1208  swap$ empty$

1209 or

1210  { pop$                    %$

1211 } 1212 { * 1213 }

1214  if$                %$
1215  behead.t

1216 }

1217 while$                                  %$
1218 %  duplicate$ "dehyphen.dewhitespace after" debugval          %$

1219 }

1220

1221 FUNCTION {format.names.for.sort}
1222 { 's :=
1223 #1 'nameptr :=
1224 s num.names$ 'numnames :=                  %$
1225 numnames 'namesleft :=
1226 % WHILE
1227 { namesleft #0 > }

1228 {

1229  % NOTE: order or name-parts ought to follow that used for leaditem
1230  % word-by-word alphabetization:
1231 %      s nameptr "{vv{ } }{ll{ }}{  jj{ }}{  ff{ }}" format.name$                  %$
1232  % letter-by-letter alphabetization, no attention to initials:
1233 %      s nameptr "{vv{}}{ll{}}{  jj{}}{  ff{}}" format.name$                  %$

1234

1235  % letter-by-letter alphabetization, with first initials exception:

1236

1237  s nameptr "{vv{}}{ll{}}{  jj{}}" format.name$         %$

1238

1239  % Each first-name token of the form of an initial ("L.") will be followed by a
1240  % space but other spaces will be removed.

1241

1242  % Algorithm: step through looking for space, which terminates name tokens,
1243  % adding non-spaces to an accumulator on the stack as we go.
1244  % When we find a space, look at the accumulator (a whole name token).
1245  % If it's length is 2 and 2d char is a period, we've got an initial, so
1246  % add the space to the accumulator.  If not, throw away the space
1247  % and step on looking for another name.

1248

1249  % FIX: would it be simpler to replace periods with spaces?
1250  %      would a period ever appear in a first-name when not an initial?
1251  %      Rev.?  Col.?  These should be alphabetized as spelled out, prob.
1252  %      but need to check \manual

1253

1254 %      duplicate$ "beginning of initials funniness; last-part is" debugval %$

1255

1256  s nameptr "{ff{ } }" format.name$ 't :=      %$
1257  % t is "F1 F2 F3 ... FN " or empty

1258

1259 %      t "initial t" debugval

1260

1261  t empty$       %$
1262  'skip$       %$
1263  { delimiter.name-part *
1264  ""        % accumulator
1265  % WHILE
1266  { t "" = not }
1267  { t head
1268 %              duplicate$ "stepping. t is [" t * "] and head is" * debugval %$
1269  duplicate$ empty$
1270  % last char will always be a space
1271  % accumulator contains whole name token
1272  { pop$       %$  throw space away for now; top of stack is now accumulator
1273 %                  duplicate$ "found a space; accumulator is" debugval %$
1274  duplicate$ %$
1275  duplicate$ text.length$ #2 =        %    2 chars long
1276  swap$ behead head period =          %$   2d char is period

1277 and

1278  % name token in accumulator is an initial, so add a space to it
1279  { space *
1280 %                      duplicate$ "found an initial" debugval %$

1281 }

1282  % name token in accumulator is not an initial
1283  'skip$      %$
1284  if$        %$
1285  % in either case, concatenate the accumulator the name on the
1286  % stack and initialize it again
1287 %                  duplicate$ "adding to name" debugval %$

1288 * "" 1289 }

1290  % accumulator contains partial name token
1291  % add t's head (not a space) to accumulator

1292 { *

1294  if$         %$
1295  behead.t

1296 }

1297  while$            %$
1298  % we always end up with an empty string on the stack
1299  pop$     %$
1300 %          duplicate$ "formatted first-name is" debugval %$
1301 %          s nameptr "{ff}{vv}{ll}{jj}" format.name$ "it came from" debugval %$

1302 }

1303  if$      %$
1304 %      duplicate$ "end of initials funniness; stack has" debugval             %$

1305

1306  format.dehyphen
1307  't :=
1308  % t now contains one name of a namelist
1309  nameptr #1 =

1310 { t 1311 }

1312  { delimiter.name *
1313  nameptr numnames =
1314  t "others" =

1315 and

1316  { "zzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzz" *

1317 } 1318 { t * 1319 }

1320  if$                  %$

1321 }

1322  if$                  %$
1323  nameptr #1 + 'nameptr :=
1324  namesleft #1 - 'namesleft :=

1325 }

1326 while$                  %$

1327 }

FIX: should be able to share code with make.leaditem.with.tags We don’t bother to format.sortify.clip, just do that once before assignment

to the ‘’ sort.key$.

1328 FUNCTION {make.principal-field.for.sort}  % FIX: note: leaditem must already be set
1329 {   principal-type author.ptype =
1330  principal-type editor.ptype =

1331 or

1332  { principal-field format.names.for.sort

1333 }

1334  { principal-type emergency.ptype =
1335  { principal-field

1336 }

1337  % organization.ptype, key.ptype, including both key.special cases
1338  % leaditem is the same as principal-field except in the key.special
1339  % cases, when we do want the leaditem not the principal-field
1340 { leaditem
1341 % letter-by-letter alphabetization:
1342  format.dehyphen.dewhitespace

1344  if$           %$

1345 }

1346 if$            %$

1347 }

4.14 Begin and end entries

begin.entry Begin an entry.

1348 FUNCTION {begin.entry}
1349 { output.begin
1350 output.bibitem

1351 }

finish.entry Conclude an entry. \stackstringnull

1352 FUNCTION {finish.entry}
1353 { add.period$                      %$
1354 output.end
1355 newline$                         %$
1356 make.annotation write$           %$

1357 }

4.15 Entry types

4.15.1 Article

article An article from a journal or magazine. Required: ‘’ author, ‘’ title, ‘’ journal, ‘’

year. Optional: ‘’ volume, ‘’ number, ‘’ pages, ‘’ month, ‘’ note. Title quoted.

FIX: this is not true any more. hmm, which warnings to give? Warnings given for no ‘’ pages; no ‘’ volume or ‘’ number; and ‘’ number but no ‘’ volume. It’s a drag, but the latter warning should remain there. The reason is that this bibstyle will produce a very misleading entry if you mistakenly omit the volume and give a number for a journal.

FIX: this is the old def of article, following output of title through end of crossref conditional.

       volume field.or.null
 %    \end{macrocode}
 %  One item (the volume) is on the stack now. Each of the four cases in the
 %  following conditionals should leave a second item on the stack.
 %    \begin{macrocode}
       number empty$                   %$
         { volume empty$                   %$
             { "empty volume and number in " cite$ * warning$
               ""
             }
             { colon
             }
           if$                   %$
         }
         { volume empty$                   %$
             { "number but no volume in " cite$ * warning$
 %    \end{macrocode}

   % FIX when a journal has no volume but a number, there should be a comma after
   % the journal, a space and then ``no. n:ppp''.  But we have already said
   % ``space output'' after the journal name, no doubt.  So we have to make this
   % decision earlier.
   %    \begin{macrocode}
                 "\unskip, no." tie * number * colon *
               }
               { "(" number * ")" * tie.or.space.connect
   %    \end{macrocode}
   % FIX: I don't think this is handling the case without pages right -- there
   % will be a colon with nothing following.  I could save either colon or
   % colon+space in a variable, and slap it on before the pages if there are pages
   % and leave it off if there are no pages.  Hmm, no: see \manaul 15.215
   % for how to handle missing pages.  Not easy.
   %
   % Now there's just one item (the volume) on the stack, so leave the following
   % as the second item.
   %    \begin{macrocode}
                 colon space *
               }
             if$
           }
         if$ *
         make.pages "pages" check.empty *
         space output
         month parenthesize space output

A crossref from an article says “See this, pages xxx.” See Chicago Manual 16:107 for this peculiar bit about the spacing after the colon.

1358 FUNCTION {article}
1359 { begin.entry
1360 make.leaditem.with.tags space output
1361 new.block
1362 make.year.nomonth "year" check.empty space output
1363 new.block
1364 make.title.inquotes "title" check.empty space output
1365 new.block
1366 crossref empty$                              %$
1367  { journal italicize "journal" check.empty space output
1368  series empty$                              %$
1369  'skip$                              %$

FIX: take this out when we finish make.vol.series.

1370  { "Achicago can not yet handle nonempty series! in " cite$ * warning$

1371 }

1372  if$                              %$
1373  volume empty$ number empty$ month empty$ pages empty$ and and and
1374  { "empty volume, number, month, and pages in " cite$ * warning$

1375 }

1376  'skip$                              %$
1377  if$                              %$
  See definition of make.vol.series.num.month.pages for why we set u in that

function.

1378  make.vol.series.num.month.pages u output

1379 }

Unusual case where you are probably citing one journal issue and numerous articles within it, so you have one entry for the issue itself, and others referring to it. Perhaps there are other cases, but I think what’s right here is to give the crossref and only the pages.

1380  { "In " make.crossref * space output
1381  make.pages "pages" check.empty comma output

1382 }

1383 if$                              %$
1384 new.block
1385 note space output
1386 finish.entry

1387 }

4.15.2 Book

book
1388 FUNCTION {book}
1389 { begin.entry
1390 make.leaditem.with.tags comma output
1391 new.block
1392 make.year.or.oyear.month "year" check.empty comma output
1393 new.block
1394 make.title.ital "title" check.empty comma output
1395 new.block
1396 make.edition comma output
1397 new.block
1398 crossref empty$                              %$
1399  { volume empty$                              %$
1400  { output.series.number

1401 }

1402  { output.volume.series

1403 }

1404  if$                              %$
1405  editor empty$ author empty$ or
1406  'skip$                              %$
1407  { new.block
1408  make.editors.secondary comma output

1409 }

1410  if$                  %$
1411  new.block
1412  make.translators comma output
1413  new.block
1414  make.illustrators comma output
1415  new.block
1416  oyear empty$                              %$
1417  { address comma output
1418  publisher "publisher" check.empty colon output

1419 }

1420  { oaddress comma output
1421  opublisher "opublisher" check.empty colon output
1422  new.block

1423  "Reprint" comma output
1424  address field.or.null comma output
1425  publisher "publisher" check.empty
1426  address empty$                              %$
1427  { comma

1428 }

1429  { colon

1430 }

1431  if$                  %$
1432  output
1433  year "year" check.empty comma output

1434 }

1435  if$                              %$

1436 }

1437  { volume empty$                              %$
1438  { "In "

1439 }

1440  { "Volume" volume tie.or.space.connect
1441  " of " *

1442 }

1443  if$                      %$
1444  make.crossref * comma output

1445 }

1446 if$                              %$
1447 new.block
1448 yearcomp empty$                              %$
1449  'skip$                              %$
1450  { "Composed in " yearcomp *
1451  comma output
1452  new.block

1453 }

1454 if$                              %$

The ‘’ fyear must be nonempty to get any information about the first-published edition,

The ‘’ fpublisher, ‘’ faddress, and ‘’ flanguage can be empty in any combination. They are all ignored if ‘’ fyear is empty, except for possibly ‘’ flanguage, which will be given with the translator if ‘’ translator is nonempty.

1455 fyear empty$
1456  'skip$
1457  { "First published "
1458  translator empty$ flanguage empty$ not and
1459  { "in " * flanguage * space *

1460 }

1461  'skip$

1462 if$

If ‘’ ftitle is empty, there will be a harmless? extra space. FIX: surely i can fix that...

1463  make.as.ftitle.ital * space output
1464  faddress empty$ fpublisher empty$ not and
1465  { "by "

1466 }

1467  { "in "

1468 }

1469  if$                  %$
1470  space output
1471  faddress space output
1472  fpublisher
1473  faddress empty$                  %$
1474  { space

1475 }

1476  { colon

1477 }

1478  if$                  %$
1479  output
1480  fyear
1481  faddress empty$ fpublisher empty$ and
1482  { space

1483 }

1484  { comma

1485 }

1486  if$ output                  %$
1487  new.block

1488 }

1489 if$                                        %$
1490 note comma output
1491 finish.entry

1492 }

4.15.3 Inbook

inbook Chicago Manual doesn’t seem to recognize parts of a book without their own titles. We make up our own version, moving the chapter and page information from the end of the ‘’ booktitle to the beginning.

1493 FUNCTION {inbook}
1494 { begin.entry
1495 make.leaditem.with.tags comma output
1496 new.block
1497 make.year.or.oyear.month "year" check.empty comma output
1498 new.block

We store an extra copy of make.chapter.of in the variableu here, so that we know whether to put pages or not after the ‘’ title sentence (with editors, and so on).

1499 crossref empty$            %$
1500  { make.chapter.of
1501  duplicate$ 'u :=              %$
1502  duplicate$ empty$
1503  { pop$                  %$
1504  pages empty$                  %$
1505  { "empty chapter, type, and pages in " cite$ * warning$
1506  "In "

1507 }

1508  { "Pages " make.pages * " in " *

1509 }

1510  if$                  %$

1511 }

1512  'skip$                  %$

1513  if$                  %$
1514  comma output
1515  make.title.ital "title" check.empty space output
1516  make.bookauthors.secondary comma output
1517  make.editors.secondary comma output
1518  volume empty$                  %$
1519  { output.series.number

1520 }

1521  { output.volume.series

1522 }

1523  if$                  %$
1524  make.edition comma output

Here we branch again on the contents of make.chapter.of. If it is not empty, then we go ahead and make.pages. If it is empty we don’t, because this means that if ‘’ pages were given, they have already been put earlier.

1525  u empty$                  %$
1526  'skip$                  %$
1527  { make.pages comma output

1528 }

1529  if$                  %$
1530  new.block
1531  make.translators comma output
1532  new.block
1533  make.illustrators comma output
1534  new.block
1535  oyear empty$                  %$
1536  { address comma output
1537  publisher "publisher" check.empty colon output

1538 }

1539  { oaddress comma output
1540  opublisher "opublisher" check.empty colon output
1541  new.block
1542  "Reprint" comma output
1543  address field.or.null comma output
1544  publisher "publisher" check.empty
1545  address empty$                  %$
1546  { comma

1547 }

1548  { colon

1549 }

1550  if$                  %$
1551  output
1552  year "year" check.empty comma output

1553 }

1554  if$                  %$

1555 }

1556  % crossref not empty
1557  { "In " make.crossref * comma output
1558  make.chapter comma output
1559  make.pages comma output
1560  chapter empty$ pages empty$ and
1561  { "empty chapter and pages in " cite$ * warning$

1562 }

1563  'skip$                  %$
1564  if$                  %$

1565 }

1566 if$                  %$
1567 new.block
1568 note comma output
1569 finish.entry

1570 }

4.15.4 Incollection

incollection
1571 FUNCTION {incollection}
1572 { begin.entry
1573 make.leaditem.with.tags comma output
1574 new.block
1575 make.year.or.oyear.month "year" check.empty space output
1576 new.block
1577 % FIX: use \field{type} to say what kind of title this is? e.g., play, poem,
1578 % chapter, story?
1579 make.title.inquotes "title" check.empty space output
1580 new.block
1581 crossref empty$                  %$
1582  { make.chapter.of duplicate$ empty$
1583  { pop$                  %$
1584  "In "

1585 }

1586  'skip$                        %$
1587  if$                  %$
1588  space output
1589  make.booktitle.ital "booktitle" check.empty space output
1590  make.bookauthors.secondary comma output
1591  make.editors.secondary comma output
1592  volume empty$                  %$
1593  { output.series.number

1594 }

1595  { output.volume.series

1596 }

1597  if$                  %$
1598  make.edition comma output
1599 % FIX: do I need this?  I don't for copleston:cusa
1600 %      make.year.or.oyear.month "year" check.empty comma output
1601  make.pages comma output
1602  chapter empty$ type empty$ and pages empty$ and                  %$
1603  { "empty chapter, type, and pages in " cite$ * warning$

1604 }

1605  'skip$                          %$
1606  if$                  %$
1607  new.block
1608  make.translators space output
1609  new.block
1610  make.illustrators space output
1611  new.block
1612  oyear empty$                  %$

1613  { address space output
1614  publisher "publisher" check.empty colon output

1615 }

1616  { oaddress space output
1617  opublisher "opublisher" check.empty colon output
1618  new.block
1619  "Reprint" comma output
1620  address comma output
1621  publisher "publisher" check.empty
1622  address empty$                  %$
1623  { comma

1624 }

1625  { colon

1626 }

1627  if$                  %$
1628  output
1629  year comma output

1630 }

1631  if$                  %$

1632 }

1633  % crossref not empty
1634  { new.block
1635  "In " make.crossref * space output
1636  make.chapter comma output
1637  make.pages comma output
1638  chapter empty$ pages empty$ and
1639  { "empty chapter and pages in " cite$ * warning$

1640 }

1641  'skip$                  %$
1642  if$                  %$

1643 }

1644 if$                  %$
1645 new.block
1646 note space output
1647 finish.entry

1648 }

4.15.5 Majorthesis

format.thesis thesis phdthesis mastersthesis majorthesis minorthesis Theses are all set in the same way, with a different type name. An entry with entry type ‘thesis’ ought to have a nonempty “” type. The other four alias entry types simply supply an explicit value for “” type, which are overridden by a nonempty “” type.

FIX The “” type field should be hooked into a variable and done with the babel package: \majorthesisname and so on.

1649 FUNCTION {format.thesis}
1650 { 'v :=
1651 begin.entry
1652 make.leaditem.with.tags comma output
1653 new.block
1654 make.year.month "year" check.empty space output
1655 new.block
1656 make.title.inquotes "title" check.empty space output

1657 new.block
1658 v "type" check.empty space output
1659 school "school" check.empty comma output
1660 address comma output
1661 new.block
1662 note space output
1663 finish.entry

1664 }

1665 FUNCTION {thesis}
1666 { type field.or.null format.thesis

1667 }

1668 FUNCTION{mastersthesis}
1669 { type empty$               %$
1670  { "Master's thesis"

1671 }

1672  { type

1673 }

1674 if$               %$
1675 format.thesis

1676 }

1677 FUNCTION {minorthesis}
1678 { mastersthesis

1679 }

1680 FUNCTION {phdthesis}
1681 { type empty$               %$
1682  { "Ph.D.~diss."

1683 }

1684  { type

1685 }

1686 if$               %$
1687 format.thesis

1688 }

1689 FUNCTION {majorthesis}
1690 { phdthesis

1691 }

4.15.6 Manual

manual Technical documentation. Required: ‘’ title (italicized). Optional: ‘’ author or ‘’ organization or ‘’ key, ‘’ address, ‘’ edition, ‘’ month, ‘’ year, ‘’ note. Logic: leaditem is first nonempty field of ‘’ author, ‘’ organization, and ‘’ key. Subsequent nonempty fields in that list are ignored. Special case: when principal-field is ‘’ organization and ‘’ key is nonempty, use ‘’ key for the label instead of ‘’ organization.

1692 FUNCTION {manual}
1693 { begin.entry
1694 make.leaditem.with.tags space output
1695 new.block
1696 make.year.month space output
1697 new.block
1698 make.title.ital "title" check.empty comma output
1699 new.block
1700 make.edition comma output

1701 new.block
1702 address space output
1703 organization colon output
1704 new.block
1705 note space output
1706 finish.entry

1707 }

4.15.7 Booklet

booklet A work that is printed and bound, but without a named publisher or sponsoring institution. Required: ‘’ title. Optional: ‘’ author, ‘’ howpublished, ‘’ address, ‘’ month, ‘’ year, ‘’ note. Title is quoted. ‘’ key used if ‘’ author is empty.

1708 FUNCTION {booklet}
1709 { begin.entry
1710 make.leaditem.with.tags comma output
1711 new.block
1712 make.year.month space output
1713 new.block
1714 make.title.inquotes "title" check.empty comma output
1715 new.block
1716 howpublished space output
1717 address comma output
1718 new.block
1719 note space output
1720 finish.entry

1721 }

4.15.8 Inproceedings

inproceedings conference An article in a conference proceedings. Required: ‘’ author, ‘’ title, ‘’ booktitle, ‘’ year. Optional: ‘’ editor, ‘’ volume or ‘’ number, ‘’ series, ‘’ pages, ‘’ address, ‘’ month, ‘’ organization, ‘’ publisher, ‘’ note. ‘’ key used where ‘’ author empty. Title quoted.

1722 FUNCTION {inproceedings}
1723 { begin.entry
1724 make.leaditem.with.tags comma output
1725 new.block
1726 make.year.month "year" check.empty space output
1727 new.block
1728 make.title.inquotes "title" check.empty space output
1729 new.block
1730 crossref empty$                  %$
1731  { make.editors.secondary space output
1732  make.booktitle.ital "booktitle" check.empty comma output
1733  volume empty$                  %$
1734  { output.series.number

1735 }

1736  { output.volume.series

1737 }

1738  if$                  %$
1739  make.edition comma output
1740  new.block

1741  organization space output
1742  address comma output
1743  publisher colon output
1744  make.pages comma output

1745 }

1746  { "In " make.crossref * space output
1747  make.chapter comma output
1748  make.pages comma output
1749  chapter empty$ pages empty$ and
1750  { "empty chapter and pages in " cite$ * warning$

1751 }

1752  'skip$                  %$
1753  if$                  %$

1754 }

1755 if$                  %$
1756 new.block
1757 note space output
1758 finish.entry

1759 }

1760 FUNCTION {conference}
1761 { inproceedings

1762 }

4.15.9 Proceedings

proceedings The proceedings of a conference. Required: ‘’ title, ‘’ year. Optional: ‘’ editor, ‘’ volume or ‘’ number, ‘’ series, ‘’ address, ‘’ month, ‘’ organization, ‘’ publisher, ‘’ note. If ‘’ editor empty, use ‘’ organization for leaditem. For label, like type manual, use nonempty ‘’ key for the label instead of ‘’ organization. Title italicized.

1763 FUNCTION {proceedings}
1764 { begin.entry
1765 make.leaditem.with.tags space output
1766 new.block
1767 make.year.month "year" check.empty space output
1768 new.block
1769 make.title.ital "title" check.empty comma output
1770 new.block
1771 make.edition comma output
1772 new.block
1773 volume empty$               %$
1774  { output.series.number

1775 }

1776  { output.volume.series

1777 }

1778 if$               %$
1779 new.block
1780 organization space output
1781 address comma output
1782 publisher colon output
1783 new.block
1784 note comma output
1785 finish.entry

4.15.10 Misc

Required: none. The rest are optional. ‘’ leaditem is ‘’ author or ‘’ key if empty. ‘’ year with ‘’ month. Then ‘’ title, quoted. Then ‘’ howpublished. Then ‘’ note.

1787 FUNCTION {misc}
1788 { begin.entry
1789 make.leaditem.with.tags comma output
1790 new.block
1791 make.year.month space output
1792 new.block
1793 title space output
1794 new.block
1795 howpublished comma output
1796 new.block
1797 note space output
1798 finish.entry

1799 }

4.15.11 Techreport

techreport A report published by a school or other institution, usually numbered within a series. Required: ‘’ author, ‘’ title, ‘’ institution, ‘’ year. Optional: ‘’ type, ‘’ number, ‘’ address, ‘’ month, ‘’ note. ‘’ key used when ‘’ author empty.

1800 FUNCTION {techreport}
1801 { begin.entry
1802 make.leaditem.with.tags comma output
1803 new.block
1804 make.year.month "year" check.empty comma output
1805 new.block
1806 make.title.inquotes "title" check.empty comma output
1807 new.block
1808 make.number.tr comma output
1809 institution "institution" check.empty comma output
1810 address comma output
1811 new.block
1812 note comma output
1813 finish.entry

1814 }

4.15.12 Unpublished

unpublished A document having an author and title, but not formally published. Required: ‘’ author, ‘’ title, ‘’ note. Optional: ‘’ month, ‘’ year. Uses ‘’ key for no ‘’ author.

1815 FUNCTION {unpublished}
1816 { begin.entry
1817 make.leaditem.with.tags comma output
1818 new.block
1819 make.year.month space output
1820 new.block
1821 make.title.inquotes "title" check.empty space output
1822 new.block

1823 note "note" check.empty space output
1824 finish.entry

1825 }

4.15.13 Default

default.type I think BibTEX is hardwired to look for this name.

1826 FUNCTION {default.type}
1827 { misc }

4.16 Sorting

chop.word \stack⟨string⟩ ⟨num⟩ ⟨string⟩⟨string⟩

% stack: S n T
%
% if S = substr(1,n,T)
%    then substr(n+1, max, T)
%    else S
%
% if T begins with S, then chop it off the beginning of T.
%

I guess we don’t bother calculating the length for efficiency’s sake.

1828 FUNCTION {chop.word}
1829 { 't :=
1830 'len :=
1831 t #1 len substring$ =               %$
1832  { t len #1 + global.max$ substring$

1833 } 1834 { t 1835 }

1836 if$               %$

1837 }

make.title.for.sort We don’t bother to format.sortify.clip, just do that once before assignment to the sort.key.

1838 FUNCTION {make.title.for.sort}
1839 { title field.or.null
1840 't :=
1841 "A " #2
1842  "An " #3
1843  "The " #4 t chop.word
1844  chop.word
1845 chop.word
1846 % letter-by-letter alphabetization:
1847 format.dehyphen.dewhitespace

1848 }

presort Set label, year.label, leaditem. Set ‘’ sort.key$ to leaditem + year.label

+ title suitably massaged.

Note that year.label.tag is not set yet; it will be set in stages in

forward.pass.add.bcds and reverse.pass.add.as.

1849 FUNCTION {presort}
1850 { set.principal-field.principal-type

1851

1852 %  "principals set for " cite$ * debugmsg        %$
1853 make.leaditem
1854 'leaditem :=

1855

1856 %  "leaditem set for " cite$ * debugmsg       %$
1857 make.label
1858 'label :=

1859

1860 %  "label set for " cite$ * debugmsg            %$
1861 oyear empty$                   %$
1862  { year

1863 }

1864  { oyear

1865 }

1866 if$                   %$
1867 % FIX: substring or text.prefix?  if purify removes special chars then no diff.
1868 field.or.null purify$ #-1 #4 substring$
1869 'year.label :=

1870 }

1871

1872 FUNCTION {sortkey<-label+year+principal+title}
1873 { label
1874 delimiter.sub-sortkey *
1875 year.label *
1876 delimiter.sub-sortkey *
1877 make.principal-field.for.sort *
1878 delimiter.sub-sortkey *
1879 make.title.for.sort *
1880 format.sortify.clip
1881 'sort.key$ :=                  %$

1882 }

1883

1884 FUNCTION {sortkey<-principal+year+year-tag}
1885 { make.principal-field.for.sort
1886 delimiter.sub-sortkey *
1887 year.label *
1888 % FIX if year.label is ever going to be a variable length, need some padding here?
1889 year.label.tag *
1890 format.sortify.clip
1891 'sort.key$ :=                  %$

1892 }

1893

1894 FUNCTION {presort.&.sortkey<-label+year+principal+title}
1895 { presort
1896 %  "presort done for " cite$ * debugmsg      %$
1897 sortkey<-label+year+principal+title
1898 %  "sortkey set to label+year+principal+title for " cite$ * debugmsg       %$

1899 }

initialize.sorting-gvars
1900 FUNCTION {initialize.sorting-gvars}

1901 { unidentified 'last.leaditem :=
1902 unidentified 'last.label.and.year :=
1903 "" 'following.year.label.tag :=
1904 #0 'last.year.tag.num :=

1905 }

5 Disambiguating citations and sorting the bibliography

The initial calculation of citations can result in ambiguous labels for some entries. Two kinds of ambiguities can result:

same principal These are disambiguated by adding a unique lowercase letter to the end of the citations for each entry in the ambiguous set: (Smith 1980a) and (Smith 1980b); (Smith et al. 1970a), (Smith et al. 1970b). The order of the letters corresonds to the order of their appearance in the bibliography (see below).

different principal Different principals can have be (initially) given the same label in two cases: (1) the principals include 4 or more names and the initial names share a surname, and (2) the principals include the same number n < 4 of names and both the initial names and each of the subsequent names when present share a surname. In each case, the entries must also have the same year+tag, that is, the same year including any lowercase letters that might have been added. In both cases, ambiguous labels are disambiguated by adding first initials to those names which differ in the first name: (P. Smith 1980), (M. Smith 1980); (Green, Z. Black, and Meyer 1980), (Green, B. Black, and Meyer 1980); (Benson, Brown, Gray, White, B. Black, and Griscom 1900), (Benson, H. Brown, Gray, White, Z. Black, and Griscom 1900). When even first initials aren’t enough to disambiguate, full names are used: (John D. Smith 1990), (John Q. Smith 1990); etc.

FIX: current behavior disambiguates all ambiguous citations by means of the lowercase letter. “A, B, C, and D 1999” together with “A, B, C, and E 1999” are going to resolve to “A et al 1999a” and 1999b. This nicely avoids disambiguation, but not what the Manual specifies.

FIX: current behavior is going to disambiguate citations that are ambiguous because authors share last but differ in first names by means of the lowercase letter: “Aa 1999” and “Ab 1999” become “Aa 1999a” and “Ab 1999b” (where ‘A’ stands for last name, and the ‘a’ right after it is the unmentioned first name). The Manual, however, specifies (16.3) disambiguating by using first initials in the label, such as “M. Brown 1999”.

add.bcds: work on getting the right year.label.tag. Compare current entry’s label plus year.label to the last entry’s, which is stored in the global string variable last.label.and.year. If they are the same, increment global integer variable last.year.tag.num and set year.label.tag to the nth lowercase letter, where n is last.year.tag.num. If they are not the

same, set last.label.and.year to the current label plus year.label, set
year.label.tag to empty string, and set last.year.tag.num to 0.

After add.bcds pass, year.label.tag will be blank for all first entries of an ambiguous series, and “b,” “c,” etc. for second entries, third entries, and so on. The “a”’s will be added in the right places by the reverse.pass.add.as function.

Handle duplicate leaditems. Compare current entry’s leaditem with the last entry’s, which is stored in the global string variable last.leaditem. When they are the same, wrap the current entry’s leaditem in \SCduplicate.

1906 FUNCTION {forward.pass.wrap.dupe.leaditems}
1907 { leaditem
1908 duplicate$ last.leaditem =                  %$
1909  % we add the brace pair at the end because this forces add.period$ to add a  %$
1910  % period after it.  FIX: if we ever want \SCdup to actually print its
1911  % argument, we must get it to print its argument minus any final period.
1912  { "\SCduplicate{" swap$ * "{}}" * 'leaditem :=                  %$

1913 }

1914  { 'last.leaditem :=

1915 }

1916 if$                  %$

1917 }

1918 FUNCTION {forward.pass.add.bcds}
1919 { label year.label * format.clip
1920 duplicate$ last.label.and.year =                  %$
1921  % same as previous entry
1922  { pop$                  %$
1923  last.year.tag.num #1 + 'last.year.tag.num :=
1924  last.year.tag.num int.to.chr$ 'year.label.tag :=                  %$

1925 }

1926  % different from previous entry
1927  { "a" chr.to.int$ 'last.year.tag.num :=                  %$
1928  "" 'year.label.tag :=
1929  'last.label.and.year :=

1930 }

1931 if$                  %$

1932 }

1933 % FIX: use a minimum of entry-variables to hold 'last' things.
1934 % FIX: change `last' to 'prev' where appropriate.  what about vars that might
1935 % be used going forward and back?  always use `prev'?

1936

reverse.pass.add.as Pass through the entries in reverse order. Set current year.label.tag to “a” on entries that precede ones with “b”s. Set following.year.label.tag to the

  current year.label.tag.
1937 FUNCTION {reverse.pass.add.as}
1938 { following.year.label.tag "b" =
1939  { "a" 'year.label.tag :=

1940 }

1941  'skip$                  %$
1942 if$                  %$
1943 year.label.tag 'following.year.label.tag :=

1944 }

begin.bib end.bib These two functions are called to begin and end the entire bibliography.

1945 FUNCTION {begin.bib}
1946 { preamble$ empty$

1947  'skip$                  %$
1948  { preamble$ write$ newline$                  %$

1949 }

1950 if$                  %$
1951 "\begin{thebibliography}{}" write$ newline$

1952 }

1953 FUNCTION {end.bib}
1954 { newline$                  %$
1955 "\end{thebibliography}" write$ newline$

1956 }

5.1 Actions

This is the main subroutine which calls all the others.

1957 READ

1958

1959 EXECUTE {init.scalars}

1960

1961 ITERATE {presort.&.sortkey<-label+year+principal+title}
1962 SORT

1963

1964 EXECUTE {initialize.sorting-gvars}

1965

1966 % add in "b"'s etc.
1967 ITERATE {forward.pass.add.bcds}
1968 % add in "a"'s
1969 REVERSE {reverse.pass.add.as}

1970

1971 ITERATE{sortkey<-principal+year+year-tag}
1972 SORT

1973

1974 EXECUTE {initialize.sorting-gvars}
1975 ITERATE {forward.pass.wrap.dupe.leaditems}

1976 1977

1978 EXECUTE {begin.bib}
1979 ITERATE {call.type$}                  %$
1980 EXECUTE {end.bib}

References

Abbey, Edward. 1968. Desert Solitaire. Ballantine.

Abbey, Edward, and Philip Hyde. 1987. Slickrock. Peregrine Smith.

Abrams, M. H. 1953. The Mirror and the Lamp: Romantic Theory and the Critical Tradition. London: Oxford University Press.

. 1973. Natural Supernaturalism: Tradition and Revolution in Romantic Literature. New York: W. W. Norton and Co.

Abrams, M. H., et al., eds. 1979. The Norton Anthology of English Literature. 4th ed. 2 vols. New York: W. W. Norton and Co.

Acheson, James. 1979. “Murphy’s Metaphysics.” Journal of Beckett Studies, no. 5:9–23.

Ackerley, C. J. 1998a. Demented Particulars: The Annotated Murphy. Published as volume 7 nos. 1–2 of the Journal of Beckett Studies.

Ackerley, Chris. 1997. “ ‘Do Not Despair’: Samuel Beckett and Robert Greene.” Journal of Beckett Studies, n.s., 6 (1): 119–24.

. 1998b. “Samuel Beckett’s Sibilants; or, Why Does Murphy Hiss?” Journal of Beckett Studies, n.s., 8 (1): 119–20.

Admussen, Richard L. 1979. The Samuel Beckett Manuscripts: A Study. Boston: G. K. Hall and Co.

Alighieri, Dante. 1982–86. The Divine Comedy. Bantam Classic ed. The Divine Comedy. Translated from the Italian by Allen Mandelbaum. Illustrated by Barry Moser. New York: Bantam Books.

. 1954a. The Inferno. Translated from the Italian by John Ciardi. New York: New American Library.

. 1954b. The Inferno. Translated from the Italian by John Ciardi. Pen-Guin.

. 1961. The Purgatorio. Translated from the Italian by John Ciardi. New York: New American Library.

. 1970. The Paradiso. Translated from the Italian by John Ciardi. New York: New American Library.

. 1982. Purgatorio: A Verse Translation. Bantam Classic ed. Volume 1 of Alighieri 86. Notes by Allen Mandelbaum and Gabriel Marruzzo, with Larry Magnus.

. 1984. Purgatorio: A Verse Translation. Bantam Classic ed. Volume 2 of Alighieri 86. Notes by Laury Magnus, Allen Mandelbaum, and Anthony Oldcorn, with Daniel Feldman.

. 1986. Purgatorio: A Verse Translation. Bantam Classic ed. Volume 3 of Alighieri 86. Notes by Anthony Oldcorn and Daniel Feldman, with Giuseppe Di Scipio.

Allison, Alexander W., et al., eds. 1983. The Norton Anthology of Poetry. 3d ed. New York and London: W. W. Norton and Co.

Amiran, Eyal. 1993. Wandering and Home: Beckett’s Metaphysical Narrative. University Park: Pennsylvania State University Press.

Andonian, Cathleen Culotta, ed. 1998. The Critical Response to Samuel Beckett. Critical Responses in Arts and Letters no. 30. Westport CT: Greenwood Press.

Anspaugh, Kelly. 1996. “ ‘Faith, Hope, and—What Was It?’: Beckett Reading Joyce Reading Dante.” Journal of Beckett Studies, n.s., 5 (1–2): 19–38.

Aquinas, Thomas, Saint. 1964. Christian Theology (Ia. I). Volume 1 of Summa Theologiae. 60 vols. Cambridge: Blackfriars.

Argyros, Alex. 1991. A Blessed Rage for Order. Studies in Literature and Science. Ann Arbor: University of Michigan Press.

Aristophanes. 1962. Four Comedies. Edited by Dudley Fitts. San Diego: Harcourt Brace Jovanovich.

Armstrong, Gordon. 1993. “Cultural Politics and the Irish Theatre: Samuel Beckett and the New Biology.” Theatre Research International 18 (3): 215– 21 (autumn).

Atanasijevic, Ksenija. 1923. The Metaphysical and Geometrical Doctrine of Bruno As Given in His Work De Triplici Minimo. Modern Concepts of Philosophy. Translated by George Vid Tomashevich. Reprint, Saint Louis, Missouri: Warren H. Green, 1972.

Auster, Paul. 1990. The New York Trilogy. Penguin Books.

Ayer, A. J., et al. 1979. Perception and Identity: Essays Presented to A. J. Ayer, With His Replies. Edited by Graham Macdonald. Cornell University Press.

Baars, Donald. 1983. The Colorado Plateau: A Geologic History. University of New Mexico.

Bair, Deirdre. 1990. Samuel Beckett: A Biography. New York: Summit Books. First published in 1978 by Harcourt Brace Jovanovich (New York) and Jonathan Cape, but this edition does not appear, technically, to be a reprint.

Baker, Phil. 1997. Beckett and the Mythology of Psychoanalysis. New York: St. Martin’s Press.

Barale, Michèle Aina, and Rubin Rabinovitz, eds. 1988. A KWIC Concordance to Samuel Beckett’s Trilogy: Molloy, Malone Dies, and The Unnamable. 2 vols. Volume 753 of Garland Reference Library of the Humanities. New York: Garland Publishing.

, eds. 1990. A KWIC Concordance to Samuel Beckett’s Murphy. Volume 1079 of Garland Reference Library of the Humanities. New York: Garland Publishing.

Bate, Walter Jackson, ed. 1970. Criticism: The Major Texts. Enl. ed. San Diego: Harcourt Brace Jovanovich.

Beckett, Samuel. 1955–58. Three Novels. New York: Grove Press.

. 1931. Proust. Reprint, Grove Press, 1957.

. c. 1934. “Echo’s Bones.” Typescript story in the Special Collections of Dartmouth College Library, Hanover, New Hampshire. 28 pages.

. 1937. “Human Wishes.” In Beckett 1984c, 155–66.

. 1949. “Three Dialogues with George Duthuit.” In Beckett 1984c, 138–45.

. 1953. Watt. Paris: Olympia Press. Reprint, New York: Grove Press, 1959.

. 1954. Waiting for Godot. First Evergreen ed. Translated from the French by the author. New York: Grove Press. Reprint, New York: Grove Press, 1956. Composed in 1948. First published as En attendant Godot in 1952. Premiered in French in 1953, in English in 1955.

. 1955. Molloy. Translated by Patrick Bowles and Samuel Beckett. New York: Grove Press.

. 1956a, 7 April. Letter to Barney Rosset of Grove Press in the Special Collections of Syracuse University Library, Syracuse, New York.

. 1956b. Malone Dies. New York: Grove Press. . 1956c. “Yellow.” In New World Writing, Volume 10 of New American

Library Mentor Books, 108–19. New York: New American Library. Also in Beckett 1972, 158–74, with minor differences.

. 1957. Murphy. New York: Grove Press. First published in 1938. . 1958a. Endgame. New York: Grove Press. . 1958b. Mexican Poetry: An Anthology. Edited by Octavio Paz. Bloom-

ington: Indiana University Press. Reprint, New York: Grove, 1985.

. 1958c. Stories and Texts for Nothing. Reprint, New York: Grove Press, 1967.

. 1958d. The Unnamable. New York: Grove Press. . 1964. How It Is. Translated from the French by the author. New York:

Grove Press. First published as Comment c’est in Paris: Les Editions de Minuit, 1961.

. 1972. More Pricks Than Kicks. New York: Grove Press. First published by Chatto and Windus, 1934.

. 1974a. First Love and Other Shorts. New York: Grove Press. . 1974b. Mercier and Camier. New York: Grove Press. . 1976. Fizzles. New York: Grove Press. . 1977. Collected Poems in English and French. New York: Grove Press. . 1980. Company. New York: Grove Press. . 1981. Ill Seen Ill Said. New York: Grove Press. . 1983a. “Letter [of 22 April 1958] from Samuel Beckett concerning

manuscript of story ‘Killachter Meadow.’ ” Review of Contemporary Fiction 3 (1): 156–57.

. 1983b. Nohow On. Riverrun Press. . 1983c. A Samuel Beckett Reader. Edited by John Calder. London: Pan

Books.

. 1983d. Worstward Ho. New York: Grove Press. . 1984a. “All That Fall.” In Beckett 1984b, 9–41. . 1984b. Collected Shorter Plays. New York: Grove Press.

. 1984c. Disjecta: Miscellaneous Writings and a Dramatic Fragment. Edited by Ruby Cohn. New York: Grove Press.

. 1984d. “Krapp’s Last Tape.” In Beckett 1984b, 53–63.

. 1984e. “That Time.” In Beckett 1984b, 225–35.

. 1984f. Three Plays: Ohio Impromptu, Catastrophe, What Where. New York: Grove Press.

. 1986. Happy Days: The Production Notebook of Samuel Beckett. Edited by James Knowlson. New York: Grove Press. First published in London: Faber and Faber, 1985.

. 1987. As the Story Was Told. Edited by Christopher B. Ricks. Cambridge, England: Rampant Lions Press.

. 1988. Stirrings Still. Illustrated by Louis le Brocquy. Blue Moon Books. Reprint, New York: North Star Line, 1991.

. 1990. As the Story Was Told: Uncollected and Late Prose. New York: Riverrun Press.

. 1992a. Endgame: With a Revised Text. Volume 2 of The Theatrical Notebooks of Samuel Beckett. Edited by S. E. Gontarski. London: Faber and Faber.

. 1992b. Krapp’s Last Tape: With a Revised Text. Volume 3 of The Theatrical Notebooks of Samuel Beckett. Edited by James Knowlson. New York: Grove Press.

. 1993a. Dream of Fair to middling Women. Edited by Eoin O’Brien and Edith Fournier. New York: Arcade Publishing in association with Riverrun Press. Composed in 1932.

. 1993b. Samuel Beckett’s Company/ Compagnie and A Piece of Monologue/ Solo: A Bilingual Variorum Edition. Volume 1400 of Garland Reference Library of the Humanities. Edited by Charles Krance. New York: Garland Publishing.

. 1994. Waiting for Godot: With a Revised Text. Volume 1 of The Theatrical Notebooks of Samuel Beckett. Edited by Dougald McMillan and James Knowlson. New York: Grove Press. First published in Great Britain: Faber and Faber, 1993.

. 1995a. Eleuthéria: A Play in Three Acts. Translated from the French by Michael Brodsky. New York: Foxrock.

. 1995b. “Ping.” In Beckett 1995c, 193–96.

. 1995c. Samuel Beckett: The Complete Short Prose, 1929–1989. Edited by S. E. Gontarski. New York: Grove Press.

. 1996. Samuel Beckett’s Mal vu mal dit/ Ill Seen Ill Said: A Bilingual, Evolutionary, and Synoptic Variorum Edition. Volume 1266 of Garland Reference Library of the Humanities. Edited by Charles Krance. New York: Garland Publishing.

. 1997. Waiting for Godot. New York: Grove Press.

. 1999a. Beckett’s Dream Notebook. Edited by John Pilling. Reading, England: Beckett International Foundation.

. 1999b. The Shorter Plays: With Revised Texts. Volume 4 of The Theatrical Notebooks of Samuel Beckett. Edited by S. E. Gontarski. New York: Grove Press. First published in London: Faber and Faber, 1999.

. 2000. Beckett in Black and Red: The Translations for Nancy Cunard’s Negro (1934). Edited by Alan Warren Friedman. Lexington: University Press of Kentucky.

Beckett, Samuel, and Alan Schneider. 1998. No Author Better Served: The Correspondence of Samuel Beckett and Alan Schneider. Edited by Maurice Harmon. Cambridge: Harvard University Press.

Beer, Ann. 1985. “Watt, Knott, and Beckett’s Bilingualism.” Journal of Beckett Studies, no. 10:37–75.

Beja, Morris, S. E. Gontarski, and Pierre Astier, eds. 1983. Samuel Beckett: Humanistic Perspectives. Columbus: Ohio State University Press.

Benét, William Rose. 1965. The Reader’s Encyclopedia. 2d ed. New York: Thomas Y. Crowell.

Benson, Michael. 1984. “Moving Bodies in Hardy and Beckett.” Essays in Criticism 34 (3): 229–43 (July).

Ben-Zvi, Linda. 1984. “Fritz Mauthner for Company.” Journal of Beckett Studies, no. 9:65–88.

Bergin, Thomas Goddard, and Max Harold Fisch. 1970. The New Science of Giambattista Vico. Cornell University Press.

Blake, William. 1966. Blake’s Job: William Blake’s Illustrations of the Book of Job. Edited by S. Foster Damon. University Press of New England.

Blisset, William. 1989. “The Syntax of Violence.” In Matthias 1989, 193–208.

Bloom, Harold. 1973. The Anxiety of Influence: A Theory of Poetry. New York: Oxford University Press.

. 1985a. “Introduction.” In Bloom 1985b, 1–5.

, ed. 1985b. Samuel Beckett. Modern Critical Views. New York: Chelsea House Publishers.

. 1988a. “Introduction.” In Bloom 1988b, 1–12.

, ed. 1988b. Samuel Beckett’s Molloy, Malone Dies, The Unnamable. Modern Critical Interpretations. New York: Chelsea House Publishers.

Boden, Margaret A., ed. 1990. The Philosophy of Artificial Intelligence. Oxford University Press.

Bomberg, David. 1988. David Bomberg. London: Tate Gallery. For the exhibition of 17 Feb to 8 May 1988.

Bonnerot, Louis. 1973–1974. “David Jones and the Notion of Fragments.” Agenda 11 (4)–12 (1):76–89.

Bonola, Roberto. 1912. Non-Euclidean Geometry: A Critical and Historical Study of its Development. Translated by H. S. Carslaw. Chicago: Open Court Publishing. Authorized translation.

Borges, Jorge Luis. 1963. Ficciones. Edited by Anthony Kerrigan. Translated from the Spanish by the editor et al. New York: Grove Weidenfeld. First published in Buenos Aires: Emecé Editores, 1956.

Boulger, James D. 1980. The Calvinist Temper in English Poetry. The Hague: Mouton Publishers.

Brater, Enoch. 1983a. “The Company Beckett Keep: The Shape of Memory and One Fabulist’s Decay of Lying.” In Beja, Gontarski, and Astier 1983, 157–71.

. 1983b. “Mis-Takes, Mathematical and Otherwise, in The Lost Ones.” Modern Fiction Studies 29 (1): 93–109 (spring).

, ed. 1986. Beckett at 80/Beckett in Context. Oxford: Oxford University Press.

. 1989. Why Beckett. New York: Thames and Hudson.

Breuer, Rolf. 1993. “Paradox in Beckett.” Modern Language Review 88 (July): 559–80.

Britannica, Encyclopaedia. 1993. Encyclopaedia Britannica. 15th ed. Chicago: University of Chicago.

Brooke, Rupert. 1968. The Letters of Rupert Brooke. Edited by Geoffrey Keynes. New York: Harcourt, Brace and World.

Bruno, Giordano. 1950. “Concerning the Cause, Principle, and One.” In The Infinite in Giordano Bruno, With a Translation of His Dialogue Concerning the Cause, Principle, and One. Translated by Sidney Greenburg. New York: King’s Crown Press.

Bryden, Mary. 1992. “Figures of Golgotha: Beckett’s Pinioned People.” In Pilling and Bryden 1992, 45–62.

. 1995. “Beckett and the Three Dantean Smiles.” Journal of Beckett Studies, n.s., 4 (2): 29–33.

. 1998. Samuel Beckett and the Idea of God. New York: St. Martin’s Press.

. 1999. “The Carved Rats of Saint-Sernin.” The Beckett Circle: Newsletter of the Samuel Beckett Society 21 (2): 9 (fall).

Bulfinch, Thomas. 1979. Bulfinch’s Mythology. New York: Gramercy Books.

Buning, Marius. 1990. “Samuel Beckett’s Negative Way: Intimations of the Via Negativa in his Late Plays.” In Jasper and Crowder 1990, Chapter 9, 129–42.

Burnet, John. 1964. Greek Philosophy: Thales to Plato. 1st ed., reset and reprinted. New York: St. Martin’s Press. First published in 1914.

Burrows, Rachel. 1989. “Interview with Rachel Burrows, Dublin, Bloomsday, 1982.” Journal of Beckett Studies, no. 11–12:6–15. Interviewed by S. E. Gontarski, Martha Feshenfeld, and Dougald McMillan.

Butler, Lance St John, ed. 1993. Critical Essays on Samuel Beckett. Critical Thought Series no. 4. Hants, England: Scolar Press.

. 1999. “Review of Samuel Beckett and the Idea of God by Mary Bryden (New York: St. Martin’s, 1998).” The Beckett Circle: Newsletter of the Samuel Beckett Society 21 (2): 10–11 (Fall).

Calder, John, ed. 1967. Beckett at 60: A Festschrift. London: Calder and Boyars.

. 1979. “Review of Samuel Beckett: A Biography by Deirdre Bair (Jonathan Cape, 1978).” Journal of Beckett Studies, no. 4:74–79.

. 1983. “Introduction.” In Beckett 1983c, 9–42.

Campbell, Joseph. 1983. Historical Atlas of World Mythology. London: Summerfield Press.

Carey, Phyllis. 1992. “Stephen Dedalus, Belacqua Shuah, and Dante’s Pietà.” In Carey and Jewinski 1992b, Chapter 8, 104–16.

Carey, Phyllis, and Ed Jewinski. 1992a. “Introduction.” In Carey and Jewinski 1992b, xi–xviii.

, eds. 1992b. Re: Joyce’n Beckett. New York: Fordham University Press.

Carnap, Rudolf. 1966. Philosophical Foundations of Physics. Edited by Martin Gardner. Basic Books.

Carroll, John S. 1906. Prisoners of Hope: an Exposition of Dante’s Purgatorio. London: Hodder and Stoughton.

Carroll, Lewis. n.d. The Complete Works of Lewis Carroll. Random House.

Cassedy, Steven. 1988. “Mathematics, Relationism, and the Rise of Modern Literary Aesthetics.” Journal of the History of Ideas 49 (1): 109–32.

Castillo, Debra A. 1982. “Beckett’s Metaphorical Towns.” Modern Fiction Studies 28 (2): 189–200 (summer).

Cheng, Vincent J. 1987. “Stephen Dedalus and the Black Panther Vampire.” James Joyce Quarterly 24 (2): 161–76 (winter).

Coe, Richard N. 1983. “Beckett’s English.” In Beja, Gontarski, and Astier 1983, 36–57.

Coetzee, J. M. 1972. “The Manuscript Revisions of Beckett’s Watt.” Journal of Modern Literature 2:472–80.

Cohen, Joseph. 1975. Journey to the Trenches: The Life of Isaac Rosenberg 1890–1918. New York: Basic Books.

Cohn, Ruby. 1970. “The Laughter of Sad Sam Beckett.” In Friedman 1970, 185–97.

, ed. 1975. Samuel Beckett: A Collection of Criticism. New York: McGraw-Hill.

. 1980. Just Play: Beckett’s Theater. Princeton: Princeton University Press.

Common Prayer. 1789. The Book of Common Prayer and Administration of the Sacraments and Other Rites and Ceremonies of the Church. New York: Oxford University Press. “This edition … was … reproduced from a certified edition … as adopted by the General Convention of 1928 and emended by subsequent Conventions.”

Conley, Tom. 1985. “Samuel Beckett: Color, Letter, and Line.” Visible Language 19 (4): 484–98 (autumn).

Connor, Steven. 1982. “Beckett’s Animals.” Journal of Beckett Studies, no. 8:29–44.

Cook, Theodore Andrea. 1914. The Curves of Life: Being an Account of Spiral Formations and Their Application to Growth in Nature, to Science and to Art; with Special reference To the Manuscripts of Leonardo da Vinci. Reprint, New York: Dover Publications, 1979.

Copleston, Frederick. 1946. Greece and Rome. Volume 1 of A History of Philosophy. 9 vols. in 3 books. Newman Press. Reprint, New York: Image Books, 1985.

. 1950a. Augustine to Scotus. Volume 2 of A History of Philosophy. 9 vols. in 3 books. Newman Press. Reprint, New York: Image Books, 1985.

. 1950b. “The Pseudo-Dionysius.” In Copleston 1950a, Chapter 9.

. 1953a. “Nicholas of Cusa.” In Copleston 1953b, Chapter 15.

. 1953b. Ockham to Suarez. Volume 3 of A History of Philosophy. 9 vols. in 3 books. Newman Press. Reprint, New York: Image Books, 1985.

. 1959. Hobbes to Hume. Volume 5 of A History of Philosophy. 9 vols. in 3 books. Newman Press. Reprint, New York: Image Books, 1985.

. 1960a. Descartes to Leibniz. Volume 4 of A History of Philosophy. 9 vols. in 3 books. Newman Press. Reprint, New York: Image Books, 1985.

. 1960b. Wolff to Kant. Volume 6 of A History of Philosophy. 9 vols. in 3 books. Newman Press. Reprint, New York: Image Books, 1985.

. 1963a. Fichte to Nietzsche. Volume 7 of A History of Philosophy. 9 vols. in 3 books. Newman Press. Reprint, New York: Image Books, 1985.

. 1963b. “Schopenhauer.” In Copleston 1963a, Chapters 13–14.

. 1966. Bentham to Russell. Volume 8 of A History of Philosophy. 9 vols. in 3 books. Newman Press. Reprint, New York: Image Books, 1985.

. 1974a. “August Comte.” In Copleston 1974c, Chapter 5.

. 1974b. “Henri Bergson.” In Copleston 1974c, Chapters 9–10.

. 1974c. Maine de Biran to Sartre. Volume 9 of A History of Philosophy. 9 vols. in 3 books. Newman Press. Reprint, New York: Image Books, 1985.

Cork, Richard. 1988. “Bomberg’s Odyssey.” In Bomberg 1988. For the exhibition

of 17 Feb to 8 May 1988. Cousineau, Thomas J. 1979. “Watt: Language as Interdiction and Consolation.”

Journal of Beckett Studies, no. 4:1–13. Crane, Gregory R., ed. 2000, March. The Perseus Project. http://www.perseus.

tufts.edu.

Cronin, Anthony. 1996. Samuel Beckett: The Last Modernist. London: Harper-

Collins. Culik, Hugh. 1983. “The Place of Watt in Beckett’s Development.” Modern

Fiction Studies 29 (1): 57–71 (spring).

. 1993a. “Mathematics as Metaphor: Samuel Beckett and and the Esthetics of Incompleteness.” Papers on Language & Literature 29 (2): 131–51 (spring).

. 1993b. “Neurological Disorder and the Evolution of Beckett’s Maternal Images.” In Butler 1993, 366–373.

Dakers, Caroline. 1987. The Countryside at War 1914–1918. London: Constable. Dallin, Leon, and Lynn Dallin. 1967. Folk Songster. Dubuque: Wm. C. Brown. Davidson, Gustav. 1967. A Dictionary of Angels Including the Fallen Angels.

New York: Free Press.

Davies, Paul. 1981. “Is Thermodynamic Gravity a Route to Quantum Gravity?” In Isham, Penrose, and Sciama 1981.

, ed. 1989. The New Physics. Cambridge: Cambridge University Press.

. 1994. The Ideal Real: Beckett’s Fiction and Imagination. Associated University Presses.

Day, Robert Adams. 1980. “How Stephen Wrote His Vampire Poem.” James Joyce Quarterly 17 (2): 183–98 (winter).

Dearlove, J. E. 1977. “ ‘Last Images’: Samuel Beckett’s Residual Fiction.” Journal of Modern Literature 6 (1): 104–126 (February).

Dearlove, Judith E. 1983. “ ‘Syntax Upended in Opposite Corners’: Alterations in Beckett’s Linguistic Theories.” In Beja, Gontarski, and Astier 1983, 122– 28.

Doherty, Francis. 1990. “Visions of Hell: Lowry and Beckett.” In Jasper and Crowder 1990, Chapter 8, 117–28.

Donne, John. 1980. Paradoxes and Problems. Edited by Helen Peters. Oxford: Clarendon Press.

Downes, Rackstraw. 1996. “Theater for the Free Mind and the Strong Stomach.” New York Times, 28 July, H5.

Drake, Stillman. 1973. Copernicus—Philosophy and Science: Bruno, Kepler, Galileo. Norwalk, Connecticut: Burndy Library.

Driver, Tom F. 1961. “Beckett by the Madeleine.” Columbia University Forum 4 (3): 21–25 (summer). Note: Driver “reconstructs” Beckett’s words “from notes made immediately after” their conversation (p. 22).

Dukes, Gerry. 1996. The Beckett Festival: Lincoln Center Festival 1996. Gate Theater. Program for a festival of theater perforamnces by the (Irish) Gate Theater company at the Lincoln Center in New York. Dukes wrote the notes for each play.

Eade, J. C. 1981. “The Seventh Scarf: A Note on Murphy.” Journal of Beckett Studies, no. 7:115–17.

Edson, Russell. 1973. The Clam Theater. Volume 64 of Wesleyan Poetry Program. Middletown: Wesleyan University Press.

Eiseley, Loren. 1979. The Star Thrower. San Diego: Harvest Books.

Elam, Keir. 1986. “Not I : Beckett’s Mouth and the Ars(e) Rhetorica.” In Brater 1986, 124–48.

Eliot, George. 1984. Middlemarch. New York: Random House.

Eliot, T. S. 1980. The Complete Poems and Plays 1909–1950. New York: Harcourt Brace Jovanovich.

Ellis, Reuben J. 1989. “ ‘Matrix of Surds’: Heisenberg’s Algebra in Beckett’s Murphy.” Papers on Language & Literature 25 (1): 120–23 (winter).

Ellis, Willis D. 1950. A Source Book of Gestalt Psychology. New York: Humanities Press.

Ell Mann, Richard. 1982. James Joyce. New and Rev. ed. New York: Oxford University Press. First published in 1959. THIS IS A TEST ENTRY ellmanntest2.

Ellmann, Richard. 1982a. James Joyce. New and Rev. ed. New York: Oxford University Press. First published in 1959.

Ellmann, Richard, Robert O’Clair, et al., eds. 1980. Norton Anthology of Modern Poetry. 2d ed. New York: W. W. Norton and Co. THIS IS A TEST ENTRY norton:modern:test.

, eds. 1988. Norton Anthology of Modern Poetry. 2d ed. New York: W. W. Norton and Co.

Ellmann, Romulus. 1980. James Joyce. New and Rev. ed. New York: Oxford University Press. First published in 1959. THIS IS A TEST ENTRY ellmanntestr.

Ell Mann, Romulus. 1980. James Joyce. New and Rev. ed. New York: Oxford University Press. First published in 1959. THIS IS A TEST ENTRY ellmanntestr2.

Ellmann, Romulus. 1982b. James Joyce. New and Rev. ed. New York: Oxford University Press. First published in 1959. THIS IS A TEST ENTRY ellmanntestrx.

. 1990. James Joyce. New and Rev. ed. New York: Oxford University Press. First published in 1959. THIS IS A TEST ENTRY ellmanntestrlate.

Ellmanna, Richard. 1982. James Joyce. New and Rev. ed. New York: Oxford University Press. First published in 1959. THIS IS A TEST ENTRY ellmanntest1:joyce.

Ell Manna, Richard. 1982. James Joyce. New and Rev. ed. New York: Oxford University Press. First published in 1959. THIS IS A TEST ENTRY ellmanntest3.

Ellmanna, Romulus. 1980. James Joyce. New and Rev. ed. New York: Oxford University Press. First published in 1959. THIS IS A TEST ENTRY ellmanntestr1.

Ell Manna, Romulus. 1980. James Joyce. New and Rev. ed. New York: Oxford University Press. First published in 1959. THIS IS A TEST ENTRY ellmanntestr3.

Empson, William. 1947. Seven Types of Ambiguity. New York: New Directions.

Esslin, Martin. 1986a. “Dionysos’ Dianoetic Laugh.” In John Calder (Publishers) 1986, 15–23.

. 1986b. “Samuel Beckett—Infinity, Eternity.” In Brater 1986, 110–23.

Euclid. 1952. The Thirteen Books of Euclid’s Elements. Volume 11 of Great Books of the Western World. Translated by Sir Thomas L. Heath. Chicago: Encylopaedia Britannica.

Farrow, Anthony. 1991. Early Beckett: Art and Allusion in More Pricks Than Kicks and Murphy. Troy, New York: Whitston Publishing.

Federman, Raymond. 1965. Journey to Chaos: Samuel Beckett’s Early Fiction. Berkeley: University of California Press.

Federman, Raymond, and John Fletcher. 1970. Samuel Beckett: His Works and His Critics; An Essay in Bibliography. Berkeley: University of California Press.

Fergusson, Francis. 1953. Dante’s Drama of the Mind: A Modern Reading of the Purgatorio. Princeton University Press.

Festugière, A. J. 1950. La Révélation d’Hermès Trismégiste. Paris. Cited in Yates 1964: page references are to Yates.

Fletcher, John. 1967. Samuel Beckett’s Art. London: Chatto and Windus.

Flexner, Stuart Berg, et al., eds. 1987. The Random House Dictionary of the English Language. 2d ed. New York: Random House.

Fowlie, Wallace. 1985. “Dante and Beckett.” In Dante Among the Moderns, edited by Stuart Y. McDougal, 128–52. Chapel Hill and London: University of North Carolina Press.

Frazer, James George. 1922. The Golden Bough: A Study in Magic and Religion. 1 vol. Abridged ed. Reprint, New York: Macmillan, 1948.

Freud, Sigmund. 1924–50. Miscellaneous Papers, 1888–1938. Volume 5 of Collected Papers. Edited by James Strachey. New York: Basic Books in arrangement with Hogarth Press and the Institute for Psycho-Analysis, London. Authorized translation from the German under the supervision of Joan Riviere.

. 1910. “ ‘The Antithetical Sense of Primal Words’: A Review of a Pamphlet by Karl Abel, Über den Gegensinn der Urworte, 1884.” In Freud 1959, Chapter 10, 184–91. Authorized translation from the German under the supervision of Joan Riviere.

. 1915. “Thoughts for the Times on War and Death.” In Freud 1959, Chapter 17, 288–317. Authorized translation from the German under the supervision of Joan Riviere.

. 1919. “Psycho-analysis and War Neuroses.” In Freud 50, Chapter 8, 83–87. Authorized translation from the German under the supervision of Joan Riviere.

. 1920. “Beyond the Pleasure Principle.” In Freud 1961b, 18:1–64.

. 1925a. “Negation.” In Freud 50, Chapter 16, 181–85. Authorized translation from the German under the supervision of Joan Riviere.

. 1925b. “A Note upon the ‘Mystic Writing-Pad.’ ” In Freud 50, Chapter 15, 175–80. Authorized translation from the German under the supervision of Joan Riviere.

. 1932. “Why War?” In Freud 50, 273–87. Authorized translation from the German under the supervision of Joan Riviere.

. 1959. Papers on Metapsychology, Papers on Applied Psycho-Analysis. Volume 4 of Collected Papers. New York: Basic Books in arrangement with Hogarth Press and the Institute for Psycho-Analysis, London. First published in 1925. Authorized translation from the German under the supervision of Joan Riviere.

. 1961a. Beyond the Pleasure Principle. Edited by James Strachey. Translated from the German by James Strachey. New York: Liveright. Reprint, New York: W. W. Norton and Co., 1990. First published in 1920. On p. xxxvii: “The present translation is a corrected reprint of the Standard Edition version” (i.e., English translation of 1955).

. 1961b. The Standard Edition of the Complete Psychological Works of Sigmund Freud. Edited by James Strachey. Translated by Joan Riviere. London: Hogarth Press and the Institute for Psycho-Analysis.

. 1989a. Civilization and Its Discontents. Reprint of the Standard Edition. Translated by James Strachey. New York: W. W. Norton and Co.

. 1989b. The Ego and the Id. Reprint of the Standard Edition. Edited by James Strachey. Translated by Joan Riviere. New York: W. W. Norton and Co.

Friebert, Stuart, and David Young, eds. 1989. The Longman Anthology of Con-

temporary American Poetry: 1950 to the Present. 2d ed. Longman. Friedman, Melvin J., ed. 1970. Samuel Beckett Now. Chicago: University of

Chicago Press.

. 1995. “Samuel Beckett: Tradition and Innovation.” Contemporary Literature 36 (2): 350–61 (summer). Review of Wandering and Home by Eyal Amiran, The Beckett Studies Reader by S. E. Gontarski, Innovation in Samuel Beckett’s Fiction by Rubin Rabinovitz, and Beckett’s Dying Words by Christopher Ricks.

Frye, Northrop. 1951. “The Archetypes of Literature.” In Bate 1970, 601–9. . 1957. “Anatomy of Criticism.” In Bate 1970, 609–25. . 1993. “The Nightmare Life in Death.” In Butler 1993, 101–108.

Gardner, Edmund G. 1913. Dante and the Mystics: A Study of the Mystical Aspect of the Divina Commedia and its Relations With Some of its Mediaeval Sources. Reprint, New York: Haskell House Publishers, 1968.

Gardner, Martin. 1961. The Scientific American Book of Mathematical Puzzles

& Diversions. 2d ed. New York: Simon and Schuster.

. 1979a. “Cyclic Numbers.” In Gardner 1979c, Chapter 10, 111–22.

. 1979b. “Cyclic Numbers.” In Gardner 1979c, Chapter 14, 169–181. . 1979c. Mathemetical Circus. New York: Alfred A. Knopf. , ed. 1994. Great Essays in Science. Buffalo: Prometheus Books. . 1997. The Night is Large: Collected Essays, 1938–1995. New York: St.

Martin’s Griffin. Garner, Jr., Stanton B. 1993. “ ‘Still living flesh’: Beckett, Merleau-Ponty, and

the Phenomenological Body.” Theatre Journal 45:443–60. Gascoigne, George. 1907. The Posies. Volume 1 of The Complete Works of George

Gascoigne. 2 vols. Edited by John W. Cunliffe. Grosse Pointe, Michigan: Scholarly Press. Reprint, Cambridge: Cambridge University Press, 1969.

Gates, David. 1996. “He Couldn’t Go On, He Went On.” New York Times Book

Review, 26 May, 4. Gatti, Hilary. 1989. The Renaissance Drama of Knowledge. London: Routledge. Geerling, Paul. 1981. Down the Grand Staircase. Westwater. Geroch, Robert. 1978. Gerneral Relativity From A to B. Chicago: University of

Chicago Press.

Gibaldi, Joseph. 1998. MLA Style Manual and Guide to Scholarly Publishing. 2d ed. New York: Modern Language Assocation of America.

Gierow, Karl Ragnar. 1971. “Presentation Address [for Beckett’s Nobel Prize].” Samuel Beckett, Björnstjerne Björnson, Pearl Buck [and] Ivan Bunin, Nobel Prize Library. New York: A. Gregory, 3–6. Author is permanent Secretary of the Swedish Academy.

Gifford, Don, and Robert J. Seidman. 1988. Ulysses Annotated. Berkeley: University of California Press.

Gleitman, Henry. 1986. Psychology. 2d ed. New York: W. W. Norton and Co.

Glover, Jon, and Jon Silkin, eds. 1989. The Penguin Book of First World War Prose. Penguin Books.

Gluck, Barbara Reich. 1979. Beckett and Joyce: Friendship and Fiction. Lewisburg: Bucknell University Press.

Gobar, Ash. 1968. Philosophic Foundations of Genetic Psychology and Gestalt Psychology: A Comparative Study of the Empirical Basis, Theoretical Structure, and Epistemological Groundwork of European Biological Psychology. The Hague: Maritnus Nihjogg.

Godfrey, Tony. 1998. Conceptual Art. Art & Ideas. London: Phaidon Press.

Gontarski, S. E. 1984. “Quad I & II : Beckett’s Sinister Mime(s).” Journal of Beckett Studies, no. 9:137–38.

. 1985a. “The Intent of Undoing.” In Gontarski 1985b, 1–21.

, ed. 1985b. The Intent of Undoing in Samuel Beckett’s Dramatic Texts. Bloomington: Indiana University Press.

, ed. 1986. On Beckett: Essays and Criticism. New York: Grove Press.

, ed. 1993. Beckett Studies Reader. Gainesville: University Press of Florida.

. 1995a. “Editing Beckett.” Twentieth Century Literature 41 (2): 190–207 (summer).

. 1995b. “From Unabandoned Works: Samuel Beckett’s Short Prose.” In Beckett 1995c, xi–xxxii.

Gordon, John. 1978. “Notes in Response to Michael Seidel’s Ulysses’ Black Panther Vampire.” James Joyce Quarterly 15 (3): 229–35 (spring).

. 1990. “Haines and the Black Panther.” James Joyce Quarterly 27 (3): 587–94 (spring).

Gordon, Lois. 1996. The World of Samuel Beckett, 1906–1946. New Haven: Yale University Press.

Gose, Jr., Eliot B. 1980. The Transformation Process in Joyce’s Ulysses. Toronto: University of Toronto Press.

Gove, Philip Babcock, et al., eds. 1993. Webster’s Third New International Dictionary of the English Language, Unabridged: A Merriam-Webster. Springfield, Massachusetts: Merriam-Webster.

Grandin, Temple, and Margaret M. Scariano. 1993. Emergence: Labelled Autistic. Novato, California: Arena Press.

Grant, Edward. 1981. Much Ado About Nothing: Theories of Space and Vacuum from the Middle Ages to the Scientific Revolution. Cambridge: Cambridge University Press.

Graver, Lawrence, and Raymond Federman, eds. 1979. Samuel Beckett: The Critical Heritage. London: Routledge and Kegan Paul.

Green, David D. 1996. “Beckett’s Dream: More Niente than Bel.” Journal of Beckett Studies, n.s., 5 (1–2): 67–80.

Green, David [D.]. 1994. “A Note on Augustine’s Thieves.” Journal of Beckett Studies, n.s., 3 (2): 77–78.

Greenberg, Valerie D. 1990. Transgressive Readings: The Texts of Franz Kafka and Max Planck. Ann Arbor: University of Michigan Press.

Greene, Robert. n.d. [1881–86]. “The Repentance of Robert Greene.” In The Life and Complete Works in Prose and Verse of Robert Greene, edited by Alexander B. Grosart, Huth Library, 15 vols., 12:151–88. London and Aylesbury: printed for private circulation. Reprint, New York: Russell and Russell, 1964.

Gregory, Richard L., ed. 1987. The Oxford Companion to the Mind. Oxford: Oxford University Press.

Hague, Rene, ed. 1980. Dai Greatcoat: A Self-portrait of David Jones in his letters. London: Faber and Faber.

Halliwell-Phillipps, James Orchard, ed. 1969. The Nursery Rhymes of England, Obtained Principally from Oral Tradition. 3d ed. with additions. Detroit: Singing Tree Press. First printed in London; reprinted Boston, Munroe and Francis, 1843.

Hallyn, Fernand. 1990. The Poetic Structure of the World: Copernicus and Kepler. Translated by Donald M. Leslie. New York: Zone Books. Distributed by MIT Press, Cambridge, Massachusetts.

Hamilton, Alice, and Kenneth Hamilton. 1976. Condemned to Life: The World of Samuel Beckett. William B Eerdmans.

Hammond, B. S. 1979. “Beckett and Pinter: Towards a Grammar of the Absurd.” Journal of Beckett Studies, no. 4:35–42.

Hardy, Thomas. 1924. The Dynasts: An Epic Drama of the War with Napoleon. Reprint, Macmillan and Co., 1954.

. 1976. The Complete Poems of Thomas Hardy. Edited by James Gibson. London: Macmillan Publishing.

. 1982. The Complete Poetical Works of Thomas Hardy. Edited by Samuel Lynn Hynes. New York: Oxford University Press.

Harper, George Mills. 1961. The Neoplatonism of William Blake. Chapel Hill: University of North Carolina Press.

. 1992. “Review of The Irish Beckett by John P. Harrington (Syracuse: Syracuse University Press, 1991).” Journal of Beckett Studies, n.s., 2 (1): 131–34.

Harrington, John P. 1982. “Pynchon, Beckett, and Entropy: Uses of Metaphor.” Missouri Review 5 (3): 129–38 (summer).

. 1991. The Irish Beckett. Syracuse: Syracuse University Press.

. 1992. “A Note on Malone Dies and Local Phenomena.” Journal of Beckett Studies, n.s., 1 (1–2): 141–43.

Harrison, Robert. 1968. Samuel Beckett’s Murphy. University of Georgia Monographs no. 15. Athens: University of Georgia Press.

Harvey, Lawrence E. 1970. Samuel Beckett: Poet and Critic. Princeton: Princeton University Press.

Hay, Eloise Knapp. 1982. T. S. Eliot’s Negative Way. Cambridge: Harvard University Press.

Hayden, Robert Earl. 1985. Collected Poems. Edited by Frederick Glaysher. New York: Liveright Publishing.

Heisenberg, Werner. 1958. Physics and Philosophy: The Revolution in Modern Science. Volume 19 of World Perspective Series. Reprint, New York: Harper and Row, 1962.

Henderson, Linda Dalrympe. 1983. The Fourth Dimension and Non-Euclidean Geometry in Modern Art. Princeton: Princeton University Press.

Heninger, S. K. 1974. Touches of Sweet Harmony: Pythagorean Cosmology and Renaissance Poetics. San Marino: Huntington Library.

. 1989. Sidney and Spenser: The Poet As Maker. University Park PA: Pennsylvania State University Press.

Henning, Sylvie Debevec. 1985. “The Guffaw of the Abderite: Murphy and the Democritean Universe.” Journal of Beckett Studies, no. 10:5–20.

Henschen, Folke. 1966. The Human Skull: A Cultural History. London: Thames and Hudson.

Heraclitus. 1970. The Cosmic Fragments. Edited by G. S. Kirk. Cambridge: Cambridge University Press.

Hibberd, Dominic, and John Onions, eds. 1986. Poetry of the Great War: An Anthology. New York: St. Martin’s Press.

Hill, Geoffrey. 1984. The Lords of Limit: Essays on Literature and Ideas. London: Andre Deutsch.

. 1985. Collected Poems. Reprint, London: Andre Deutsch, 1986.

. 1989. “A Pharisee to Pharisees: Reflections on Vaughan’s ‘The Night.’ ” English 38 (summer): 97–113.

Hills, Paul. 1981. “The Art of David Jones.” In Jones 1981b. Exhibition catalog.

Hobson, Harold. 1956. “Samuel Beckett, Dramatist of the Year.” International Theater Annual, no. 1:153–55.

Holman, C. Hugh, and William Harmon. 1986. A Handbook to Literature. 5th ed. New York: Macmillan. Based on the original edition by William Flint Thrall and Addison Hibbard.

Homer. 1963. The Odyssey. First Vintage Classics ed. Translated by Robert Fitzgerald. Illustrated by Jackie Schuman. Reprint, New York: Random House, 1990. Text as corrected and reset in 1961.

Hopkins, Gerard Manley. 1953. Poems and Prose. Edited by W. H. Gardner. PenGuin Books.

Horowitz, Irving Louis. 1952. The Renaissance Philosophy of Giordano Bruno. New York: Coleman-Ross.

Housman, A. E. 1896. A Shropshire Lad. Reprint, Waterville, Maine: Colby College Library, 1946.

. 1922. Last Poems. New York: Henry Holt and Co.

Hume, Anthea. 1984. Edmund Spenser: Protestant Poet. Cambridge: Cambridge University Press.

Iser, Wolfgang. 1985. “The Pattern of Negativity in Beckett’s Prose.” In Bloom 1985b, 125–36.

Isham, C. J., R. Penrose, and D. W. Sciama, eds. 1981. Quantum Gravity 2: A Second Oxford Symposium. Oxford: Clarendon Press.

Israel, Calvin. 1979. “Review of Samuel Beckett: A Biography by Deirdre Bair (New York: Harcourt Brace Jovanovich, 1978).” Journal of Beckett Studies, no. 4:80–85.

Jacobus, Mary, Evelyn Fox Keller, and Sally Shuttleworth, eds. 1990. Body/politics: Women and the Discourses of Science. New York: Routledge.

James, William. 1890. The Principles of Psychology. 2 vols. New York: Henry Holt and Co. Reprint, New York: Dover Publications, 1950.

Jasper, David, and Colin Crowder, eds. 1990. European Literature and Theology in the Twentieth Century: Ends of Time. New York: St. Martin’s Press.

Jewinski, Ed. 1992. “James Joyce and Samuel Beckett: From Epiphany to Anti-Epiphany.” In Carey and Jewinski 1992b, Chapter 11, 160–74.

Joachim, H. H. 1951. Aristotle: The Nicomachean Ethics. Edited by D. A. Rees. Oxford: Clarendon Press.

John Calder (Publishers). 1986. As No Other Dare Fail: For Samuel Beckett on his 80th Birthday. London: John Calder.

Johnson, Mark. 1987. The Body in the Mind: The Bodily Basis of Meaning, Imagination, and Reason. Chicago: University of Chicago Press.

Jolas, Eugene, ed. 1949. Transition Workshop. New York: Vanguard Press.

Jones, Anthony. 1981a. “The French Murphy: From ‘rare bird’ to ‘cancre.’ ” Journal of Beckett Studies, no. 6:37–50.

Jones, David. 1937. In Parenthesis: seinnyessit e gledyf ym penn maeu. Reprint, London: Faber and Faber, 1963. Composed in 1932.

. 1959. Epoch and Artist. New York: Chilmark Press (distributed by Random House).

. 1974. The Sleeping Lord and Other Fragments. London: Faber and Faber.

. 1993. “An Unpublished Appreciation of Gerard Manley Hopkins.” Agenda 31 (3): 72–80. Edited by Kathleen Henderson Staudt.

Jones, David Michael. 1981b. David Jones. London: Tate Gallery. Exhibition catalog.

Joyce, James. 1939. Finnegans Wake. New York: PenGuin Books.

. 1959a. The Critical Writings of James Joyce. Edited by Ellsworth Mason and Richard Mason. Reprint, Ithaca, New York: Cornell University Press, 1989.

. 1959b. “Politics and Cattle Disease.” In Joyce 1959a, Chapter 47, 238–41.

. 1961. Ulysses. First Vintage International ed. Reprint, New York: Random House, 1990. First published in 1922.

. 1968. Dubliners. Rev. ed. Viking. Reprint, New York: Penguin Books, 1976. First published in 1916 (USA).

. 1986. Ulysses: The Corrected Text. First Vintage Books ed. Edited by Hans Walter Gabler et al. New York: Random House. First published in 1922.

. 1993. A Portrait of the Artist as a Young Man. Case Studies in Contemporary Criticism. Edited by R. B. Kershner. Boston: St. Martin’s Press, Bedford Books.

Juliet, Charles. 1995. Conversations with Samuel Beckett and Bram van Velde. Translated from the French by Janey Tucker. Academic Press Leiden. First published as Rencontres avec Bram van Velde and Rencontre avec Samuel Beckett in Montpellier: Editions Fata Morgana, 1973, 1986.

Jung, C. G. 1966a. The Collected Works of C. G. Jung. Edited by Sir Herbert Read, Michael Fordham, and Gerhard Adler. Translated by R. F. C. Hull. Princeton University Press.

. 1966b. “Psychology and Literature.” In Jung 1966a, 84–105.

. 1968a. Analytical Psychology: Its Theory and Practice. The Tavistock Lectures. New York: Vintage Books.

. 1968b. “Lecture 3.” In Jung 1968a, 78–113.

Keller, Evelyn Fox. 1985. Reflections on Gender and Science. New Haven: Yale University Press.

Kelsey, Michael. 1986. Canyon Hiking Guide to the Colorado Plateau. Kelsey.

. 1987. Hiking and Exploring the Paria River. Kelsey.

Kelsey, Michael, and Dee Ann Finken. 1990. Hiking Utah’s San Rafael Swell. Kelsey.

Kennedy, Sighle. 1971. Murphy’s Bed: A Study of Real Sources and Sur-real Associations in Samuel Beckett’s First Novel. Lewisburg, Pennsylvania: Bucknell University Press.

. 1990. “Beckett’s ‘Schoolboy Copy’ of Dante: A Handbook for Liberty.” Dalhousie French Studies 19 (fall–winter): 11–19. Special issue titled “Exile and Transcendence.”

Kenner, Hugh. 1968. Samuel Beckett: A Critical Study. New ed. Berkeley: University of California Press.

. 1975. “Shades of Syntax.” In Cohn 1975, 21–31.

Kirk, G. S., J. E. Raven, and Malcolm Schofield. 1983. The Presocratic Philosophers : A Critical History with a Selection of Texts. 2d ed. Cambridge: Cambridge University Press.

Klein, A. M. 1950. “The Black Panther (A Study in Technique).” Accent: A Quarterly of New Literature 10 (2): 139–55 (winter).

Klein, Felix. 1894. The Evanston Colloquium Lectures on Mathematics. New York: Macmillan and Co.

. 1924. Elementary Mathematics From an Advanced Standpoint. 3d ed. 2 vols. Translated by E. R. Hedrick and C. A. Noble. Dover Publications.

. 1928. Development of Mathematics in the Nineteenth Century. Volume 9 of Lie Groups. Translated by M. Ackerman. Reprint, Brookline, Massachusetts: Math Sci Press, 1979.

Knowlson, James, ed. 1971. Samuel Becket: An Exhibition Held at Reading University Library, May to July 1971. London: Turret Books.

. 1983. “Beckett’s ‘Bits of Pipe.’ ” In Beja, Gontarski, and Astier 1983, 16–25.

. 1986. “Ghost Trio/Geister Trio.” In Brater 1986, 193–207.

. 1996. Damned to Fame: The Life of Samuel Beckett. New York: Simon and Schuster.

Knowlson, James, and John Pilling. 1979. Frescoes of the Skull: The Later Prose and Drama of Samuel Beckett. London: John Calder.

Knuth, Donald Ervin. 1986. The TEXbook. 16th printing, revised. Volume A of Computers & Typesetting. Illustrated by Duane Bibby. Reading MA: Addison Wesley. First published in 1984.

Koffka, K. 1935. Principles of Gestalt Psychology. New York: Harcourt, Brace and World.

Kopczyński, W., and A. Trautman. 1992. Spacetime and Gravitation. Revised translation. Translated from the Polish by Jerzy Baldyga and Antoni Pol. Warsaw: Polish Scientific Publishers. First published as Czasoprzestrzeń i Grawitacja in Warsaw: Państwowe Wydawnictwo Naukowe, 1984.

Krance, Charles. 1983. “Odd Fizzles: Beckett and the Heavenly Sciences.” Bucknell Review 27 (2): 96–107.

Kroll, Jeri L. 1977. “The Surd as Inadmissible Evidence: The Case of Attourneygeneral v. Henry McCabe.” Journal of Beckett Studies, no. 2:47–58.

. 1978. “Belacqua as Artist and Lover: ‘What a Misfortune.’ ” Journal of Beckett Studies, no. 3:10–39.

Krutch, Joseph Wood. 1952. The Desert Year. W. Sloane.

. 1971. The Voice of the Desert. W. Sloane.

Lake, Carlton, ed. 1984. No Symbols Where None Intended: A Catalogue of Books, Manuscripts, and Other Materials Relating To Samuel Beckett in the collections of the Humanities Research Center. Austin: University of Texas at Austin Humanities Research Center.

Lakoff, George. 1987. Women, Fire, and Dangerous Things: What Categories Reveal about the Mind. Chicago: University of Chicago Press.

Lakoff, George, and Mark Johnson. 1980. Metaphors We Live By. Chicago:

University of Chicago Press. Lambrechtse, Rudi. 1985. Hiking the Escalante. Wasatch. Lamport, Leslie. 1986. L^{A}TEX: A Document Preparation System. Reading MA:

Addison Wesley. Lanczos, Cornelius. 1970. Space Through the Ages: The Evolution of Geometrical

Ideas from Pythagoras to Hilbert and Einstein. London: Academic Press. Laplanche, Jean. 1976. Life and Death in Psychoanalysis. Translated by Jeffrey

Mehlman. Baltimore: Johns Hopkins University Press. Lawlor, Robert. 1982. Sacred Geometry: Philosophy and Practice. The Illustrated

Library of Sacred Imagination. New York: Crossroad Publishing. Lawrence, D. H. 1979. The Letters of D. H. Lawrence. Edited by James T.

Boulton. Cambridge University Press. Lees, Heath. 1984. “Watt, Music, Tuning and Tonality.” Journal of Beckett

Studies, no. 9:5–24. Leggett, A. J. 1987. The Problems of Physics. Oxford University Press. Lerner, Lawrence S., and Edward A. Gosselin. April 1973. “Giordano Bruno.”

Scientific American. Levy, Jay A. 1992. “Conversations with Samuel Beckett.” American Scholar 61

(1): 124–31 (winter). Lewis, Wyndham. 1980. Wyndham Lewis. London: Lund Humphries in asso-

ciation with the City of Manchester Art Galleries. Exhibition held at the Manchester City Art Galley from 1 Oct to 15 Nov 1980.

Lightman, Alan. 1992. Great Ideas in Physics. New York: McGraw Hill. Limerick, Patricia Nelson. 1985. Desert Passages: Encounters with American

Deserts. University of New Mexico. Linden, Stanton J. 1978. “Herbert and the Unveiling of Diana: Stanza Three of

‘Vanitie.’ ” George Herbert Journal 1 (spring): 30–37. Lingenfelter, Richard E., and Richard A. Dwyer, eds. 1988. Death Valley Lore:

Classic Tales of Fantasy, Adventure, and Mystery. University of Nevada Press. Lippman, Carlee. 1993. “Proprioceptive Deficit in Beckett’s Not I .” Journal of

Beckett Studies, n.s., 2 (2): 81–83. Lizhi, Fang. 1991. “Form and Physics.” Partisan Review 58 (4): 656–664. Lodge, David. 1968. “Some Ping Understood.” Encounter 30 (2): 85–89 (Febru-

ary). Also in McCarthy 1986, 119–27. Long, Joseph. 1996. “The Reading of Company: Beckett and the Bi-Textual

Work.” Forum for Modern Language Studies 32 (4): 314–28 (October). Longair, Malcolm. 1989. “The New Astrophysics.” In Davies 1989. Lorenz, Konrad. 1981. The Foundations of Ethology. Translated from the German

by Konrad Z. Lorenz and Robert Warren Kickert. New York: Springer-Verlag. Madison, Gary Brent. 1981. The Phenomenology of Merleau-Ponty. Athens:

Ohio University Press. Cited in Garner 1993, 451.

Mann, Thomas. 1936. Death in Venice and Seven Other Stories. Translated by H. T. Lowe-Porter. Reprint, Vintage Books, 1989.

Maor, Eli. 1987. To Infinity and Beyond: A Cultural History of the Infinite. Boston: Birkhäuser.

Marculescu, Ileana. 1989. “Beckett and the Temptation of Solipsism: ‘Esse est aut percipere aut percipi.’ ” Journal of Beckett Studies, no. 11–12:53–64.

Matthias, John, ed. 1989. David Jones: Man and Poet. Orono, Maine: National Poetry Foundation [at?] University of Maine.

McCarthy, Patrick A., ed. 1986. Critical Essays on Samuel Beckett. Critical Essays on Modern British Literature. Boston: G. K. Hall and Co.

. 1998. “Review of In Principle, Beckett is Joyce, ed. Friedhelm Rathjen.” Journal of Beckett Studies, n.s., 8 (1): 141–45.

McHugh, Roland. 1991. Annotations to Finnegans Wake. Rev. ed. Baltimore: Johns Hopkins University Press.

McLaughlin, Peter. 1990. Kant’s Critique of Teleology in Biological Explanation: Antimony and Teleology. Lewiston: E. Mellen Press.

McMillan, Dougald. 1975. “Samuel Beckett and the Visual Arts: The Embarrassment of Allegory.” In Cohn 1975, 121–35.

Melville, Herman. 1855? Moby-Dick. On the internet.

. 1988. Moby-Dick, or The Whale. Volume 6 of The Writings of Herman Melville. 15 vols. Edited by Harrison Hayford, Hershel Parker, and G. Thomas Tanselle. Evanston, Illinois: Northwestern University Press and The Newberry Library.

Mercier, Vivian. 1986. “Poet and Mathematician.” Hermathena, no. 141 (winter): 66–71.

Merwin, W. S. 1970. The Carrier of Ladders. New York: Atheneum.

Michel, Paul Henri. 1973. The Cosmology of Giordano Bruno. Translated by R. E. W. Maddison. Ithaca, New York: Cornell University Press.

Milton, John. Paradise Lost. Project Gutenberg digital edition. Contact

<hart@vmd.cso.uiuc.edu>.

. 1952–55. Paradise Lost. Volume 1 of Poetical Works. 2 vols. Edited by Helen Darbishire. Oxford: Clarendon Press. First published in 1667.

. 1968. Paradise Lost and Paradise Regained. Signet Classic Poetry Series. Edited by Christopher Ricks. PenGuin Books.

. 1998a. “Paradise Lost.” In Milton 1998b, 296–710.

. 1998b. The Riverside Milton. Edited by Roy Flannagan. Boston: Houghton Mifflin.

Minihan, George, and Aidan Higgins. 1996. Samuel Beckett: Photographs. New York: George Braziller. Photographs by Minihan, text by Higgins.

Miskinis, Steven. 1996. “Enduring Recurrence: Samuel Beckett’s Nihilistic Poetics.” ELH 63 (4): 1047–1067 (winter).

Mitchell, Breon. 1986. “A Beckett Bibliography: New Works 1976–1982.” Modern Fiction Studies 29 (1): 131–52 (spring).

Montgomery, Angela. 1991. “Beckett and Science: Watt and the Quantum Universe.” Comparative Criticism 13:171–81.

Mood, John J. 1971. “ ‘The Personal System’—Samuel Beckett’s Watt.” Publications of the Modern Language Association of America 86 (2): 255–65.

Moore, A. W. 1990. The Infinite. London: Routledge.

Morot-Sir, Edouard. 1988. “Grammatical Insincerity in The Unnamable.” In Bloom 1988b, 131–44.

Morris, William, ed. 1981. The American Heritage Dictionary of the English Language. Boston: Houghton Mifflin.

Murphy, P. J., Werner Huber, Rolf Breuer, and Konrad Schoell. 1994. Critique of Beckett Criticism: A Guide to Research in English, French, and German. Literary Criticism in Perspective. Columbia: Camden House.

Nabokov, Vladimir Vladimirovich. 1980. Lectures on Literature. New York: Harcourt Brace Jovanovich.

NeXT Computer and Merriam-Webster. 1992. Webster’s Ninth New Collegiate Dictionary. First Digital ed. NeXT Computer and Merriam-Webster.

Nichols, Gary. 1986. River Runners’ Guide to Utah and Adjacent Areas. University of Utah.

Nicholson, Marjorie Hope. Mountain Gloom and Mountain Glory.

Nietzsche, Friedrich Wilhelm. 1968. The Will to Power. Edited by Walter Kaufmann. Translated by Walter Kaufmann and R. J. Hollingdale. New York: Vintage Books.

Niven, Ivan Morton, Herbert S. Zuckerman, and Hugh L. Montgomery. 1991. Theory of Numbers. 5th ed. New York: Wiley.

Nurmi, Martin K. 1972. Blake’s “Marriage of Heaven and Hell”: A Critical Study. New York: Haskell House.

O’Brien, Eoin. 1986a. The Beckett Country: An Exhibition for Samuel Beckett’s Eightieth Birthday. Dublin: Black Cat Press. catalogue by Eoin O’Brien & James Knowlson; photography by David Davidson; “this exhibition was first held in The Library, University of Reading [newline] May 1986.

. 1986b. The Beckett Country: Samuel Beckett’s Ireland. New York: Black Cat Press in association with Riverrun Press.

. 1992. “Foreward.” In Beckett 1993a, xi–xx.

Odum, Howard Washington, and Guy Benton Johnson. 1968. The Negro and His Songs: A Study of Typical Negro Songs in the South. New York: Greenwood Publishing, Negro Universities Press. First published in 1925. Reprint of edition of 1925 by University of North Carolina Press, Chapel Hill.

O’Hara, J. D. 1981. “Jung and the Narratives of Molloy.” Journal of Beckett Studies, no. 7:19–47.

. 1992. “Freud and the Narrative of Moran.” Journal of Beckett Studies, n.s., 2 (1): 47–63.

. 1994. “Beckett Backs Down: From Home to Murphy via Valéry.” Journal of Beckett Studies, n.s., 3 (2): 37–55.

. 1997. Samuel Beckett’s Hidden Drives: Structural Uses of Depth Psychology. Gainesville: University Press of Florida.

O’Malley, Kevin. 1984. “Review of Theatre Workbook 1, Samuel Beckett: Krapp’s Last Tape, ed. James Knowlson.” Journal of Beckett Studies, no. 9:145–50.

O’Neill, Kevin C. 1992. “Two Trilogies: Notes on Beckett and Dante.” Romance Notes 32 (3): 235–40 (spring).

Opie, Iona, and Peter Opie, eds. 1952. The Oxford Dictionary of Nursery Rhymes. Corrected ed. Oxford: Clarendon Press. First published in 1951.

Oppenheim, Lois, and Marius Buning, eds. 1996. Beckett On and On …. Madison: Fairleigh Dickinson University Press.

Orr, M. A. 1956. Dante and the Early Astronomers. FIX NEED ORIGINAL PUBLISHER AND ADDRESS. Reprint, Port Washington, New York: Kennikat Press, 1969. First published in 1913.

Ovid (Publius Ovidius Naso). 1977. Ovid in Six Volumes. 3d ed. Loeb Classical Library. Translated by Frank Justus Miller. Cambridge: Harvard University Press. Composed in 0007.

Peacham, Henry. 1954. The Garden of Eloquence. Gainesville: Scholars’ Facsimiles and Reprints.

Penrose, Roger. 1981. “Time-asymmetry and Quantum Gravity.” In Isham, Penrose, and Sciama 1981.

. 1986. “Gravity and state vector reduction.” In Quantum Concepts in Space and Time, edited by R. Penrose and C. J. Isham. Oxford: Clarendon Press.

. 1989. The Emporer’s New Mind: Concerning Computers, Minds, and The Laws of Physics. New York: Oxford University Press.

Perkins, David. Is Literary History Possible?

, ed. 1967. English Romantic Writers. San Diego: Harcourt Brace Jovanovich.

Pilling, John. 1976. Samuel Beckett. London: Routledge and Kegan Paul.

, ed. 1994. The Cambridge Companion to Beckett. Cambridge: Cambridge University Press.

. 1996. “A Short Statement with Long Shadows: Watt’s Arsene and His Kind(s).” In Oppenheim and Buning 1996, 61–68.

. 1998. Beckett Before Godot. Cambridge: Cambridge University Press.

Pilling, John, and Mary Bryden, eds. 1992. The Ideal Core of the Onion: Reading Beckett Archives. Beckett International Foundation.

Pingree, David, ed. 1986. Picatrix, the Latin Version of the Ghayat al-hakim. London: Warburg Institute. Authorship attributed to Maslamah ibn Ahmad al-Majriti.

Pompa, Leon. 1976. In Tagliacozzo and Verene 1976, 125–40.

Porter Abbott, H. 1977. “A Grammar for Being Elsewhere.” Journal of Modern Literature 6 (1): 39–46 (February).

. 1983. “The Harpooned Notebook: Malone Dies and the Conventions of Intercalated Narrative.” In Beja, Gontarski, and Astier 1983, 71–79.

Posnock, Ross. 1981. “Beckett, Valéry and Watt.” Journal of Beckett Studies, no. 6:51–62.

Powell, John Wesley. 1961. The Exploration of the Colorado River and Its Canyons. Dover.

Preminger, Alex, and T. V. F. Brogan, eds. 1993. The New Princeton Encyclopedia of Poetry and Poetics. Princeton: Princeton University Press.

Putnam, Samuel, Maida Castelhun Darton, George Reavey, and J. Bronowski, eds. 1931. France, Spain, England and Ireland. Volume 1 of The European Caravan: An Anthology of the New Spirit in European Literature. Brewer, Warren and Putnam. Subsequent volumes were not published.

Rabinovitz, Rubin. 1983. “Unreliable Narrative in Murphy.” In Beja, Gontarski, and Astier 1983, 58–70.

. 1984. The Development of Samuel Beckett’s Fiction. Urbana: University of Illinois Press.

. 1986. “Style and Obscurity in Samuel Beckett’s Early Fiction.” In McCarthy 1986, 21–29.

. 1989. “Beckett and Psychology.” Journal of Beckett Studies, no. 11– 12:65–77.

. 1990. “Using the Concordance: Murphy and Repetition.” In Barale and Rabinovitz 1990, xi–xxix.

. 1991. “Samuel Beckett, Revisionist Translator.” Review 13:273–81. Review of Beckett and Babel: An Investigation into the Status of the Bilingual Work by Brian Fitch, Toronto: University of Toronto Press, 1988.

. 1992. Innovation in Samuel Beckett’s Fiction. Urbana: University of Illinois Press.

. 1995. “Samuel Beckett’s Revised Aphorisms.” Contemporary Literature 36 (2): 203–25.

Rawlins, Ian. 1953. Aesthetics and the Gestalt. Edinburgh: Thomas Nelson and Sons.

Read, Herbert. 1955a. In Read 1955b, 3–35.

. 1955b. The Philosophy of Modern Art. New York: Meridian Books.

Redfern, Walter. 1998. “A Funny-bone to Pick with Beckett.” Journal of Beckett Studies, n.s., 8 (1): 101–19.

Redmond, James, ed. 1990. Drama and Philosophy. Themes in Drama no. 12. Cambridge: Cambridge University Press.

Renaud, Madeleine. 1967. “Beckett the Magnificant.” In Calder 1967, 81–83.

Reynolds, Mary T. 1981. Joyce and Dante: The Shaping Imagination. Princeton: Princeton University Press.

Richardson, Kenneth, ed. 1969. Twentieth Century Writing: A Reader’s Guide to Contemporary Literature. London: Newnes Books.

Ricks, Christopher B. 1993. Beckett’s Dying Words. The Clarendon Lectures, 1990. New York: Oxford University Press.

Ricoeur, Paul. 1970. Freud and Philosophy. The Terry Lectures. Translated by Denis Savage. New Haven: Yale University Press.

Riding, Alan. 1996. “Samuel Beckett Gets an Irish Embrace.” New York Times, 28 July, H5, H12.

Riquelme, John Paul. 1983. Teller and Tale in Joyce’s fiction: Oscillating Perspectives. Baltimore: Johns Hopkins University Press.

Robinson, James E. 1990. “Bruno and Beckett: Coincidence of Contraries.” In Redmond 1990, 171–83.

Robinson, Michael. 1979. “From Purgatory to Inferno: Beckett and Dante revisited.” Journal of Beckett Studies, no. 5:69–82.

Rose, Lloyd. 1994. “Way to Godot! Washington Shakespeare’s Boffo Beckett.” Washington Post, 25 May, C11.

Rosenberg, Isaac. 1979. The Collected Works of Isaac Rosenberg: Poetry, Prose, Letters, Paintings, and Drawings. Edited by Ian Parsons. New York: Oxford University Press.

Rotman, Brian. About Nothing: The Semiotics of Zero.

Rucker, Rudy. Infinity and the Mind.

Rudin, Walter. 1976. Principles of Mathematical Analysis. 3d ed. International Series in Pure and Applied Mathematics. New York: McGraw-Hill. First published in 1953.

Sacks, Oliver W. 1985. The Man Who Mistook His Wife for a Hat and Other Clinical Tales. New York: Summit Books.

Salmon, Merrilee H., John Earman, Clark Glymour, James G. Lennox, Peter Machamer, J. E. McGuire, John D. Norton, Wesley C. Salmon, and Kenneth F. Schaffner. 1992. Introduction to the Philosophy of Science. Englewood Cliffs, New Jersey: Prentice Hall.

Schopenhauer, Arthur. 1883. The World as Will and Idea. 3 vols. Translated from the German by R. B. Haldane and J. Kemp. Reprint, London: Kegan Paul, Trench, Trübner and Co., 1896.

. 1958. The World as Will and Representation. 2 vols. Translated from the German by E. F. J. Payne. Indian Hills, Colorado: Falcon’s Wing Press.

Sciama, D. W. 1981. “The Irreversible Thermodynamics of Black Holes.” In Isham, Penrose, and Sciama 1981.

Scruton, Roger. 1983a. The Aesthetic Understanding: Essays in the Philosophy of Art and Culture. Manchester, England: Carcanet Press.

. 1983b. “Beckett and the Cartesian Soul.” In Scruton 1983a, Chapter 16, 222–41.

. 1986. Spinoza. Oxford: Oxford University Press.

. 1999. The Aesthetics of Music. Oxford University Press.

Sebha, Gregor. 1987. The Dream of Descartes. The Journal of the History of Philosophy Monograph Series. Edited by Richard A. Watson. Carbondale: Southern Illinois University Press.

Segrè, Elisabeth Bregman. 1977. “Style and Structure in Beckett’s Ping: That Something Itself .” Journal of Modern Literature 6 (1): 127–47 (February).

Seidel, Michael. 1976. “Ulysses’ Black Panther Vampire.” James Joyce Quarterly 13 (4): 415–27 (summer).

Shakespeare, William. 1988. The Complete Works. Digital ed. Edited by Stanley Wells, Gary Taylor, et al. Oxford: Oxford University Press and NeXT Digital Press.

Sharp, Cecil J. 1932. English Folk Songs from the Southern Appalachians: Comprising 274 Songs and Ballads with 968 Tunes, Including 39 Tunes Contributed by Olive Dame Campbell. 2d and enl. ed. Edited by Maud Karpeles. London: Oxford University Press. First published in 1917.

Shaw, Harry. 1972. Concise Dictionary of Literary Terms. New York: McGraw-Hill.

Shenker, Israel. 1956. “Moody Man of Letters: A Portrait of Samuel Beckett, Author of the Puzzling Waiting for Godot.” New York Times, 6 May, sec. 2, pp. 1, 3.

Shimony, Abner. 1989. “The Conceptual Foundations of Quantum Mechanics.” In Davies 1989, 373–95.

. 1993. Search for a Naturalistic World View. Volume 2 vols. Cambridge: Cambridge University Press.

Shipley, Joseph T., ed. 1970. Dictionary of World Literary Terms, Forms, Technique, Criticism. Rev. and enl. ed. London: George Allen and Unwin.

Shlain, Leonard. 1991. Art & Physics: Parallel Visions in Space, Time, and Light. New York: Quill.

Sidney, Sir Philip. 1970. “An Apology for Poetry.” In Bate 1970, 82–106.

Silkin, Jon, ed. 1981. The Penguin Book of First World War Poetry. 2d ed. London: Penguin Books.

Simpson, J. A., and E. S. C. Weiner, eds. 1989. The Oxford English Dictionary. 2d ed. New York: Oxford University Press.

Singer, Dorothea Waley. 1968. Giordano Bruno: His Life and Thought (With Annotated Translation of His Work On the Infinite Universe and Worlds). New York: Greenwood Press.

Singer, Irving. 1984. The Nature of Love. Chicago: University of Chicago Press.

. 1992. Meaning in Life : The Creation of Value. New York: Free Press.

Skeat, Walter W. 1882. A Concise Etymological Dictionary of the English Language. Corrected impression. Reprint, London: Oxford University Press, 1927.

Smith, Anna. 1993. “Proceeding by Aporia: Perception and Poetic Language in Samuel Beckett’s Worstward Ho.” Journal of Beckett Studies, n.s., 3 (1): 21–37.

Smith, Sir William, and Sir John Lockwood. 1976. Chambers Murray Latin-English Dictionary. Edinburgh and London (respectively): W and R Chambers and John Murray. Also known as A Smaller Latin-English Dictionary in a 1933 edition.

Solomon, Philip H. 1975. The Life After Birth: Imagery in Samuel Beckett’s Trilogy. Romance Monographs no. 15. University, Mississippi: Romance Monographs.

Sondheimer, Ernst, and Alan Rogerson. 1981. Numbers and Infinity: A Historical Account of Mathematical Concepts. Cambridge: Cambridge University Press.

Sorabji, Richard. 1988. Matter, Space, and Motion. Ithaca: Cornell University Press.

Sorley, Charles Hamilton. 1990. The Collected Letters of Charles Hamilton Sorley. Edited by Jean Moorcroft Wilson. London: Cecil Woolf.

Stachel, John. “Changes in Concepts of Space and Time Brought about by Relativity.” Paper received from author in spring 1993. Do not know if published. 34 pages.

Stegner, Wallace. 1954. Beyond the Hundredth Meridian. Houghton Mifflin.

Steiner, George. 1996. “Leastness.” New Yorker, 16 September, 92–95.

Steppe, Wolfhard, and Hans Walter Gabler. 1985. A Handlist to James Joyce’s Ulysses: A Complete Alphabetical Index to the Critical Reading Text. Volume 582 of Garland Reference Library of the Humanities. Garland.

Stern, Richard. 1991. “Samuel Beckett.” Salmagundi, no. 90–91 (spring– summer): 179–90.

Stevens, Wallace. 1954a. Collected Poems. New York: Knopf.

. 1954b. The Collected Poems of Wallace Stevens. New York: Knopf. Reprint, New York: Vintage Books, 1990.

Stevenson, Kay Gilliland. 1986. “Belacqua in the Moon: Beckett’s Revisions of ‘Dante and the Lobster.’ ” In McCarthy 1986, 36–46.

Swift, Jonathan. 1969. A Tale of a Tub. Edited by Edward Hodnett. New York: AMS Press. First published in New York: Columbia University Press, 1930.

Swinger, Ann. 1975. Run, River, Run: A Naturalist’s Journey Down One of the Great Rivers of the American West. University of Arizona.

Tagliacozzo, Giorgio, and Donald Phillip Verene, eds. 1976. Giambattista Vico’s Science of Humanity. Johns Hopkins University Press.

Tawney, R. H. 1953. “The Attack.” In The Attack and Other Papers. New York: Harcourt Brace and Co.

Tennyson, Lord Alfred. 1971. Tennyson’s Poetry. Edited by Jr. Robert W. Hill. W. W. Norton and Co.

Topsfield, Valerie. 1988. The Humour of Samuel Beckett. New York: St. Martin’s Press.

Torretti, Roberto. 1978. Philosophy of Geometry from Riemann to Poincaré. Dordrecht, Holland: D. Reidel Publishing.

Trismegistus, Hermes. 1657. Hermes Mercurius Trismegistus his Divine pymander [microform] : in seventeen books : together with his second book called Asclepius, containing fifteen chapters with a commentary / translated formerly out of the Arabick into Greek, and thence into Latine, and Dutch, and now out of the original into English by Dr. Everard. London: “Printed by J. S. for Thomas Brewster”.

Turner, Frederick. 1992. “The School of Night.” In Natural Classicism: Essays on Literature and Science, 111–50. Charlottesville: University Press of Virginia.

Tuveson, Ernest Lee. 1982. The Avatars of Thrice Great Hermes: An Approach to Romanticism. Lewisburg: Bucknell University Press.

Unger, Peter. 1979. “I Do Not Exist.” In Ayer et al. 1979, Chapter 10, 235–51.

University of Chicago Press. 1982. The Chicago Manual of Style. 13th ed. Chicago: University of Chicago Press.

. 1993. The Chicago Manual of Style. 14th ed. Chicago: University of Chicago Press.

van Leunen, Mary-Claire. 1979. A Handbook for Scholars. Knopf.

van Velde, Bram. 1989. Bram van Velde: Musee National d’Art Moderne, Centre Georges Pompidou. Collection Classiques du XXe siecle, ISSN 0760-5153. Edited by Claire Stoullig. Paris: Le Centre. Catalog of an exhibition held 19 Oct 1989 to 1 Jan 1990.

Verdicchio, Massimo. 1989. “Exagmination Round the Fictification of Vico and Joyce.” James Joyce Quarterly 26 (4): 531–39 (summer).

Vernon, William Warren. 1907. Readings on the Purgatorio of Dante: Chiefly Based on the Commentary of Benvenuto da Imola. 3d ed., rev., 2 vols. London: Methuen. First published in 1889.

Virgil (Publius Vergilius Maro). 1909. The Aeneid. Internet Wiretap online ed. Volume 13 of Harvard Classics. Translated by John Dryden. New York: P. F. Collier and Son.

. 1972. The Aeneid of Virgil: Books 1–6. Edited by R. D. Williams. Macmillan.

. 1983. The Aeneid. Translated by Robert Fitzgerald. New York: Vintage Books.

von Goethe, Johann Wolfgang. 1961. Faust. Translated by Walter Kaufmann. New York: Doubleday. Part one and selections from part two.

. 1976. Faust. Norton Critical ed. Edited by Cyrus Hamlin. Translated by Walter Arndt. New York: W. W. Norton and Co.

Wald, Robert M. 1981. “Black Holes, Thermodynamics, and Time-reversibility.” In Isham, Penrose, and Sciama 1981.

. 1984. General Relativity. Chicago: University of Chicago Press.

Walker, D. P. 1958. Spiritual and Demonic Magic from Ficino to Campanella. University of London: Warburg Institute. Cited in Yates 1964.

Weiner, Andrew D. 1978. Chapter 1 of Sir Philip Sidney and the Poetics of Protestantism: A Study of Contexts. Minneapolis: University of Minnesota Press.

Weyl, Hermann. 1952. Symmetry. Princeton: Princeton University Press.

White, Hayden. 1976. In Tagliacozzo and Verene 1976, 65–85.

Whitman, Walt. 1965. Leaves of Grass. Edited by Sculley Bradley and Harold W. Blodgett. New York University Press. Reprint, New York: W. W. Norton and Co., 1973. First published in 1891.

Whyte, Lancelot Law. 1961. Essay on Atomism, From Democritus to 1960.

Middletown, Connecticut: Wesleyan University Press. Wilde, Oscar. 1989. Oscar Wilde. Oxford Authors. Edited by Isobel Murray.

Oxford: Oxford University Press. Will, Clifford. 1989. “The Renaissance of General Relativity.” In Davies 1989. Williams, Terry Tempest, and John Telford. 1989. Coyote’s Canyon. Gibbs

Smith. Wilson, Edward O. 1984. Biophilia. Cambridge: Harvard University Press.

. 1988. On Human Nature. Cambridge: Harvard University Press.

. 1998. Consilience: The Unity of Knowledge. New York: Knopf.

. 2000. Sociobiology: The New Synthesis. Twenty-Fifth Anniversary Edition. Cambridge, Massachusetts: Belknap Press.

Wind, Edgar. 1968. Pagan Mysteries in the Renaissance. New York: Barnes

and Noble. Winston, Mathew. 1977. “Watt’s First Footnote.” Journal of Modern Literature

6 (1): 69–82 (February). Withycombe, E. G. 1969. The Oxford Dictionary of English Christian Names.

Corrected 2d ed. Oxford: Oxford University Press. Wolfson, Harry Austryn. 1973. Studies in the History of Philosophy and Reli-

gion. 2 vols. Edited by Isadore Twersky and George H. Williams. Harvard University Press.

Wolosky, Shira. 1995. Language Mysticism: The Negative Way of Language in

Eliot, Beckett, and Celan. Stanford: Stanford University Press. Work, John Wesley. 1915. Folk Song of the American Negro. Nashville, Tennessee.

Reprint, New York: Negro Universities Press, 1969. Yates, Frances A. 1964. Giordano Bruno and the Hermetic Tradition. London:

Routledge and Kegan Paul. Zeifman, Hersh. 1975. “Religious Imagery in the Plays of Samuel Beckett.” In

Cohn 1975, 85–94.

. 1983. “Come and Go: A Criticule.” In Beja, Gontarski, and Astier 1983, 137–44.

Zurbrugg, Nicholas. 1989. “Review of Samuel Beckett Special Issue of Irish University Review: A Journal of Irish Studies 14 (1), spring 1984.” Journal of Beckett Studies, no. 11–12:163–66.

Zwinger, Ann. 1978. Wind in the Rock. Harper and Row.

Index

Numbers written in italic refer to the page where the corresponding entry is described; numbers underlined refer to the code line of the definition; numbers in roman refer to the code lines where the entry is used.

Symbols

\" . . . . . . . . . . . . . 1139 \# . . . . . . . . . . . . . 1192 \| . . . . . . . . . . . . . 1125

A

\and . . . . . . . . . . . . . \underline{101} \article . . . . . . . . \underline{1358} \author.q . . . . . . . . \underline{344}

B

\begin . . . . . . 512, 1951

\begin.bib . . . . . . \underline{1945}
\begin.entry . . . . . \underline{1348}

\bibitem . . . . . . . . . 220 \BibTeX . . . . . . . . . 1126 \blank.q . . . . . . . . . \underline{344} \book . . . . . . . . . . . \underline{1388} \booklet . . . . . . . . \underline{1708} \builtin 1106, 1187, 1190

C

\calc.labels . . . . . \underline{1003}
\check.empty . . . . . . \underline{267}

\check.q . . . . . . . . . \underline{394}

\chop.word . . . . . . \underline{1828}

\citeNP . . . . . . . . . . 907 \colon . . . . . . . . . . . \underline{124} \comma . . . . . . . . . . . \underline{124}

\conference . . . . . . \underline{1722}

D

\debugmsg . . . . . . . . \underline{79} \debugval . . . . . . . . \underline{79} \def . . . . . . . . . . 1, 8–13

\default.type . . . . \underline{1826}
\delimiter.name . . . \underline{124}
\delimiter.name-part

. . . . . . . . . . . \underline{124}

\delimiter.sub-sortkey

. . . . . . . . . . . \underline{124}

\docdate . . . . . . . . . 13

\DoXPackageS . . . . . . . 9

E

\editor.q . . . . . . . . \underline{344}

\either.or.check . . \underline{333}

\end . . . . . . . . 514, 1955 \end.bib . . . . . . . . \underline{1945} \entry . . . . . . . . . . 1048 \etal . . . . . . . . . . . . \underline{124} \evar . . . . . . . . . . . 1047

F \field . 1048, 1061, 1577

                          \field.or.null . . . . \underline{229}

\file . . . . . . . . . . . . . 1 \filedate . . . . . . . . 12 \fileinfo . . . . . . . . . 8

                          \fileversion . . . . . . 11
                          \finish.entry . . . . \underline{1352}
                          \format.clip . . . . . . \underline{255}
                          \format.n.dashify . \underline{298}
                          \format.names.firstfirst

. . . . . . . . . . . \underline{473}

                          \format.names.for.label

. . . . . . . . . . . \underline{406}

                          \format.names.for.sort

. . . . . . . . . . \underline{1105}

                          \format.names.lastfirst

. . . . . . . . . . . \underline{441}

                          \format.sortify.clip

. . . . . . . . . . . \underline{259}

                          \format.thesis . . . \underline{1649}

\function . . 1134, 1196

G \gvar . . . . . . . . . . . 1047

H \hyphen . . . . . . . . . . \underline{124}

I \inbook . . . . . . . . . \underline{1493} \inc.j . . . . . . . . . . . \underline{344}

                          \incollection . . . . \underline{1571}
                          \init.scalars . . . . . \underline{89}
                          \initelyHavECitationS

. . . . . . . . . . . . 10

                          \initialize.sorting-gvars

. . . . . . . . . . \underline{1900}

                          \inproceedings . . . \underline{1722}
                          \italicize . . . . . . . \underline{237}

K \key.q . . . . . . . . . . . \underline{344}

M

                                                   \majorthesis . . . . . \underline{1649}
                                                   \make.annotation . . \underline{508}
                                                   \make.as.ftitle.ital

. . . . . . . . . . . \underline{653}

                                                   \make.bookauthors.secondary

. . . . . . . . . . . \underline{598}

                                                   \make.booktitle.ital

. . . . . . . . . . . \underline{653}

                                                   \make.chapter . . . . . \underline{976}
                                                   \make.chapter.of . . \underline{961}
                                                   \make.crossref . . . . \underline{906}
                                                   \make.edition . . . . . \underline{518}
                                                   \make.editors.secondary

. . . . . . . . . . . \underline{583}

                                                   \make.illustrators . \underline{614}
                                                   \make.leaditem.with.tags

. . . . . . . . . . . \underline{570}

                                                   \make.number.tr . . . \underline{990}
                                                   \make.title.for.sort

. . . . . . . . . . \underline{1838}

                                                   \make.title.inquotes

. . . . . . . . . . . \underline{669}

                                                   \make.title.ital . . \underline{653}
                                                   \make.translators . \underline{614}
                                                   \make.vol.series.num.month.pages

. . . . . . . . . . . \underline{686}

                                                   \make.year.month . . \underline{531}
                                                   \make.year.nomonth . \underline{531}
                                                   \make.year.or.oyear.month

. . . . . . . . . . . \underline{552}

\manual 1194, 1252, \underline{1692}

                                                   \mastersthesis . . . \underline{1649}
                                                   \minorthesis . . . . . \underline{1649}
                                                   \multipage.p . . . . . . \underline{275}

N

                                                   \new.block . . . . . . . \underline{173}
                                                   \new.sentence . . . . . \underline{180}

\newblock . . . . . . . . 196 \not . . . . . . . . . . . . . \underline{101}

O \or . . . . . . . . . . . . . \underline{101}

                                                   \organization.q . . . \underline{344}

\output . . . . . . . . . . \underline{210}

                                                   \output.begin . . . . . \underline{166}
                                                   \output.bibitem . . . \underline{218}

\output.end . . . . . . . \underline{170}
\output.internal . . \underline{187}
\output.series.number

. . . . . . . . . . . \underline{909}

\output.volume.series

. . . . . . . . . . . \underline{939}

P

\parenthesize . . . . . \underline{246}

\period . . . . . . . . . . \underline{124}

\phdthesis . . . . . . \underline{1649}

\presort . . . . . . . . \underline{1849}

\proceedings . . . . . \underline{1763}

R

                          \reverse.pass.add.as

. . . . . . . . . . \underline{1937}

S \SCcite . . . . . . . . . . 220 \SCdup . . . . . . . . . . 1910

                          \SCduplicate . . . . . 1912
                          \set.principal-field.principal-type

. . . . . . . . . . \underline{1003}

\space . . . . . . . . . . . \underline{124}

                          \start.a.q . . . . . . . \underline{340}

T

                          \techreport . . . . . . \underline{1800}

                                                   \textitswitch . . . . . 242

\thesis . . . . . . . . . \underline{1649} \tie . . . . . . . . . . . . . \underline{124}

                                                   \tie.or.space.connect

. . . . . . . . . . . \underline{324}

U

                                                   \UnexpandableProtect

. . . . . . . . . . . 220

                                                   \unidentified . . . . . \underline{124}
                                                   \unpublished . . . . . \underline{1815}

W

                                                   \Wrapquotes . . . . . . . 674